### v1

# Low-Light Video Enhancement via Spatial-Temporal Consistent Illumination and Reflection Decomposition

Abstract
Low-Light Video Enhancement (LLVE) seeks to restore dynamic and static scenes plagued by severe invisibility and noise. One critical aspect is formulating a consistency constraint specifically for temporal-spatial illumination and appearance enhanced versions, a dimension overlooked in existing methods. In this paper, we present an innovative video Retinex-based decomposition strategy that operates without the need for explicit supervision to delineate illumination and reflectance components. We leverage dynamic cross-frame correspondences for intrinsic appearance and enforce a scene-level continuity constraint on the illumination field to yield satisfactory consistent decomposition results. To further ensure consistent decomposition, we introduce a dual-structure enhancement network featuring a novel cross-frame interaction mechanism. This mechanism can seamlessly integrate with encoder-decoder single-frame networks, incurring minimal additional parameter costs. By supervising different frames simultaneously, this network encourages them to exhibit matching decomposition features, thus achieving the desired temporal propagation. Extensive experiments are conducted on widely recognized LLVE benchmarks, covering diverse scenarios. Our framework consistently outperforms existing methods, establishing a new state-of-the-art (SOTA) performance.
1 Introduction
Low-light enhancement aims to enhance underexposed images and videos captured in low-light conditions Xu et al. (2022); Wang et al. (2021a), improving their visual quality while reducing noise. This is a longstanding and important topic in computer vision, with wide-ranging applications, such as improving portrait photography on mobile devices Ignatov et al. (2017); Hasinoff et al. (2016) and supporting various downstream tasks, including nighttime face recognition Ma et al. (2022); Wang et al. (2022a) and vehicle detection Fu et al. (2023a); Wu et al. (2022b).
Figure 1: Our method consistently achieves SOTA performance on different LLVE datasets with the same network architecture.
Numerous methods have been proposed to enhance underexposed images using supervised learning Wang et al. (2021b); Xu et al. (2022, 2023b); Cai et al. (2023); Wang et al. (2022b), yielding impressive results. However, these methods still struggle to provide satisfactory performance when applied to videos, a popular media format. The key challenge when enhancing videos is the need for consistent enhancement results across corresponding locations in different frames, which vary both spatially and temporally. Current methods rely on paired training data, where the ground truth ensures this consistency. Unfortunately, there is a scarcity of real-world, high-quality, spatially aligned video pairs for dynamic scenes, resulting in limited generalization capabilities for unseen videos with varying conditions.
One solution is to incorporate parameterized enhancement models with physical significance into the enhancement process, reducing the reliance on training data, e.g., Retinex theory Land (1977). According to this theory, an observed image $I$ can be decomposed into its illumination component $L$ and reflectance component $R$ using the equation $I=L\otimes R$, where $\otimes$ denotes element-wise multiplication. Several approaches Zhang et al. (2019); Wu et al. (2022a); Fu et al. (2023b) have been developed to predict both $R$ and $L$ under normal illumination conditions for a low-light input image. However, a suitable strategy for obtaining the Retinex decomposition in normal-light conditions for low-light input videos is currently lacking. Ensuring both temporal and spatial consistency in decomposition is crucial, yet previous strategies were primarily designed for individual images, thus posing a challenge for video processing.
In this paper, we introduce an innovative approach to attain a consistent Retinex decomposition for normal-light video outputs, ensuring consistency in both spatial and temporal dimensions. We leverage cross-frame correspondences and real-world physical continuity constraints to achieve this decomposition without the need for explicit supervision. Furthermore, to enhance the consistency of the decomposition, we’ve developed an efficient dual-structure enhancement network featuring a novel cross-frame interaction mechanism.
Differing from some previous methods Wang et al. (2019a, 2021a), which directly assume the reflection term as the enhanced version, our approach follows Cai et al. (2017); Wei et al. (2018); Wu et al. (2022a); Fu et al. (2023b), i.e., aims to recover both the reflection and illumination components of normal-light images from the provided low-light inputs. Previous methods have verified that satisfying image-level reflection and illumination decomposition can be achieved with suitable reflection and illumination priors via a data-driven manner. In this paper, we aim to design appropriate priors for video data to implement the decomposition.
We recognize that the reflectance component $R$ should represent the physical properties of objects, which should remain consistent across various observation perspectives Wei et al. (2018). To achieve this, we first calculate corresponding locations in videos and utilize these computed spatial-temporal correspondences, along with uncertainty values, to link the reflection terms across different frames. Conversely, the illumination terms are subject to a spatial smoothness loss to model real-world physical constraints Wei et al. (2018). This approach allows us to establish a consistency constraint specifically for temporal-spatial illumination and appearance in the enhanced videos, without requiring additional supervision information.
Additionally, the architecture of the enhancement networks plays a crucial role in shaping the decomposition results. It is important that different frames share features and receive simultaneous supervision, promoting the alignment of feature representations to synthesize Retinex parameters. In our paper, we introduce a novel dual network structure capable of achieving such feature propagation. Two frames are used as inputs, and their features are shared within the designed Cross-Frame Interaction Module (CFIM). CFIM facilitates interaction through a combination of long-range cross-frame attention computation and short-range channel-spatial fusion, ensuring the accurate and comprehensive merging of features from both global and local perspectives. CFIM can be seamlessly integrated into any encoder-decoder single-image enhancement structure by inserting it at the deepest hidden layer. In this paper, we illustrate this with the simplest U-Net Ronneberger et al. (2015) as an example, and it’s worth noting that CFIM comprises only 0.4M parameters.
We performed experiments on various widely recognized video enhancement datasets Wang et al. (2021a); Chen et al. (2019); Fu et al. (2023a); Zhang et al. (2021). Our results, both quantitative and qualitative, consistently demonstrate the effectiveness and state-of-the-art (SOTA) performance of our framework, as shown in Fig. 1. Furthermore, we conducted a large-scale user study involving 100 participants, which clearly showcased the superiority of our enhanced results in terms of human subjective perceptions.
In summary, our contributions are three-fold.
  • We introduce a novel canonical form for LLVE that predicts spatial-temporal consistent Retinex decomposition for normal-light outputs. This innovative decomposition strategy leverages spatial-temporal correspondences and continuity.
  • We design a new LLVE network that facilitates interaction among the features of different frames, leading to a consistent decomposition.
  • We conduct extensive experiments on public datasets, illustrating the effectiveness of our proposed framework.
2 Related Work
Low-Light Image Enhancement. To enhance the quality of a low-light video, image enhancement methods can be applied on a frame-by-frame basis. In recent years, learning-based Low-Light Image Enhancement (LLIE) techniques Yan et al. (2014, 2016); Lore et al. (2017); Cai et al. (2018); Zamir et al. (2020); Xu et al. (2020); Zeng et al. (2020); Kim et al. (2021); Zhao et al. (2021); Zheng et al. (2021); Wang et al. (2021b); Liu et al. (2021); Yang et al. (2021b); Jiang et al. (2021); Yang et al. (2021a) have made significant advancements, with a primary focus on supervised approaches thanks to the abundance of image pairs available for training. Xu et.al. Xu et al. (2022) introduced spatial-varying operations, considering Signal-to-Noise Ratio (SNR) as a prior factor. However, when applying image enhancement algorithms to individual frames, the issue of flickering often arises.
Low-Light Video Enhancement. In addition to the need for LLIE, there is a growing demand for video enhancement, considering the widespread use of videos as a popular data format on the internet and in photographic equipment. Various approaches have been proposed in this context Chen et al. (2019); Triantafyllidou et al. (2020); Ye et al. (2023); Lv et al. (2023); Xu et al. (2023a); Fu et al. (2023a). Danai et.al. Triantafyllidou et al. (2020) introduced a data synthesis mechanism to generate dynamic video pairs from SID Chen et al. (2019). Wang et.al. Wang et al. (2021a) proposed a multi-branch network that simultaneously estimates noise and illumination, suitable for videos with severe noise conditions. Liu et.al. Liu et al. (2023) and Liang et.al. Liang et al. (2023) used prior event information to learn enhancement mapping for brightening videos. Xu et.al. Xu et al. (2023a) designed a parametric 3D filter tailored for enhancing and sharpening low-light videos. Recently, Fu et.al. Fu et al. (2023a) introduced a video enhancement method called LAN, which iteratively refines illumination and adaptively adjusts it. However, it’s important to note that LAN lacks an explicit constraint for maintaining consistent reflectance and illumination decomposition.
Several datasets for video enhancement, encompassing both static Chen et al. (2019); Jiang and Zheng (2019); Wang et al. (2019b); Triantafyllidou et al. (2020) and dynamic motions Wang et al. (2021a); Fu et al. (2023a), have been introduced. In this paper, we introduce a novel LLVE method, designed to enhance effects on these datasets. Our decomposition method explicitly and consistently models reflectance and illumination for all frames, with the guidance of the videos themselves.
3 Method
We commence by revisiting the formulation of the Retinex theory in Sec. 3.1. Following that, we provide a detailed exposition of our decomposition strategy specifically tailored for video data in Sec. 3.2. Subsequently, in Sec. 3.3, we delve into the description of our cross-frame interaction. Finally, the training loss for our framework will be introduced in detail.
Figure 2: Our framework provides a comprehensive solution that includes explicit and consistent modeling for Retinex decomposition of enhanced normal-light outputs across different frames. To achieve this, we enforce consistent features in the reflectance terms across different frames by leveraging computed correspondences in the temporal dimension of videos ($\mathcal{L}(R_{t})$). Simultaneously, we ensure that the illumination terms exhibit a spatially continuous distribution ($\mathcal{L}(L_{t})$), aligning with real-world scenarios. Furthermore, our network incorporates cross-frame interaction and simultaneous supervision of different frames within a video, encouraging consistent Retinex features among these frames. For a more detailed visual representation, please refer to Fig. 3.
3.1 Retinex Model
Retinex Theory. When presented with an image $I$ (in this paper, we denote the low-light image as $I_{d}$ and the normal-light image as $I_{n}$, with $d$ and $n$ serving as subscript abbreviations), we assume its decomposition into illumination $L$ and reflectance $R$ follows the Retinex theory Land and McCann (1971); Fu et al. (2023b), as
$I=L\otimes R,$ (1)
where $\otimes$ denotes the element-wise multiplication, $L$ and $R$ can be formulated for different channels Wang et al. (2019a); Cai et al. (2023), i.e., the channel number is 3 for the sRGB domain. In this context, $L$ describes the light intensity of objects, which is expected to exhibit piece-wise continuity. On the other hand, $R$ represents the physical properties of the objects, encompassing textures and details observed in the data $I$. If we obtain $L$ and $R$ in the normal-light conditions for a given input low-light image, the target normal-light image can be acquired accordingly. Numerous methods have been proposed to address this decomposition problem Zhang et al. (2019); Wu et al. (2022a); Fu et al. (2023b), and their common objective can be summarized as
$\mathcal{L}=\|I-L\otimes R\|+\mathcal{L}(L)+\mathcal{L}(R),$ (2)
where $\mathcal{L}(L)$ and $\mathcal{L}(R)$ denote the constraints for illumination $L$ and reflectance $R$, respectively.
Shortcoming of current Retinex models for videos. Several prior approaches have been devised to attain this desired decomposition by imposing different constraints on the two components, such as structural and smoothness constraints Li et al. (2018); Wei et al. (2018). Nevertheless, there is a scarcity of satisfactory Retinex decomposition prediction tailored for sequential video data. The issue concerning the decomposition of video data can be summarized as
$\mathcal{L}_{v}=\mathbb{E}_{t=1:T}[\mathcal{L}_{R,t}+\mathcal{L}(L_{t})+\mathcal{L}(R_{t})],\quad\mathcal{L}_{R,t}=\|I_{t}-L_{t}\otimes R_{t}\|,$ (3)
where $\mathbb{E}$ is the average operation, $T$ is the number of frame. Compared with Eq. 2, we need to simultaneously guarantee the decomposition quality and the consistency among different frames’ results.
Video data guide Retinex models by themselves. We have discovered that video data itself can play a crucial role in facilitating the achievement of a satisfactory Retinex decomposition prediction for normal-light outputs, thus obviating the need for external priors typically used in low-light enhancement. By establishing correspondences among different frames, we can impose specific constraints: the reflectance of each frame, denoted as $R_{t}$, should faithfully represent the intrinsic texture of objects within the target scene, remaining consistent regardless of changes in viewpoint. Likewise, the illumination of each frame, denoted as $L_{t}$, should exhibit the desired continuity consistent with real-world physical properties. For further elaboration, please refer to Sec. 3.2.
Problem formulation. We adopt the supervised setting. Given a clip of low-light data $I_{d,t},t\in[1,T]$, there is a paired normal-light data $I_{n,t},t\in[1,T]$. We aim to directly obtain the decomposition of $I_{n,t}$ from $I_{d,t}$ using network $f$, as
$\hat{L}_{n,t},\hat{R}_{n,t}=f(I_{d,t}),\hat{I}_{n,t}=\hat{L}_{n,t}\otimes\hat{R}_{n,t},$ (4)
where $\hat{L}_{n,t}$ and $\hat{R}_{n,t}$ are the estimated targets (several examples are displayed in Fig. 4), and $\hat{I}_{n,t}$ is the predicted enhancement result.
Figure 3: The lightweight cross-frame interaction mechanism (CFIM) to propagate different frames’ features, along with the cross-frame attention and spatial-channel fusion. Cross-frame interaction can be employed in the deep feature space of arbitrary single-image encoder-decoder frameworks, and we choose U-Net here.
3.2 Spatial-Temporal Consistent Decomposition
Motivation. Video can be conceptualized as static/dynamic 3D multi-view data Wang et al. (2023b, a). Prior research has demonstrated that the multi-view data itself can be harnessed to decompose the view-dependent and view-independent elements, which respectively encapsulate the intrinsic texture and view-altering illuminations Wang et al. (2023a). Thus, once we establish correspondence relationships among $I_{d,t}$ for all $t$, we can apply the view-independent constraint to derive the reflectance component, denoted as $\hat{R}_{n,t}$, as it represents the intrinsic properties of the target scene. Subsequently, obtaining the illumination component, denoted as $\hat{L}_{n,t}$, becomes achievable through the utilization of the inherent image reconstruction loss, in conjunction with adherence to the continuity assumption.
Implementation of correspondences. Obtaining correspondences for $I_{d,t}$, $t\in[1,T]$, can be a challenging task due to the presence of various degradations, including visibility issues and noise. Fortunately, we have access to corresponding normal-light data, which significantly aids in establishing these correspondences, as illustrated in Fig. 2, where $I_{d,t}$ and $I_{n,t}$ are pixel-wise aligned.
Given two frames $I_{n,t_{1}}$ and $I_{n,t_{2}}$, correspondences can be determined using the prediction network $F$ Edstedt et al. (2023). We assume the detection of $M$ correspondences, denoted as ${c_{m}=(x_{1,m},y_{1,m},x_{2,m},y_{2,m}),m\in[1,M]}$. Each correspondence, represented by $c_{m}$, consists of four coordinates: $(x_{1,m},y_{1,m})$ represents the pixel coordinate in $I_{n,t_{1}}$, while $(x_{2,m},y_{2,m})$ signifies the coordinate in $I_{n,t_{2}}$. Furthermore, each correspondence is associated with an uncertainty value $u_{m}$, determined through a data-driven approach.
Constraints formulation. To satisfy the constraint that the corresponding pixels in the reflectance part are matched, we can formulate $\mathcal{L}(R_{t})$ as
$\small\mathcal{L}(R_{t_{1},t_{2}})=\mathbb{E}_{m\in[1,M]}u_{m}\|\hat{R}_{n,t_{1}}[x_{1,m},y_{1,m}]-\hat{R}_{n,t_{2}}[x_{2,m},y_{2,m}]\|.$ (5)
The reconstruction item in Eq. 3 can be denoted as
$\small\mathcal{L}_{R,t}=\|I_{n,t}-\hat{L}_{n,t}\otimes\hat{R}_{n,t}\|.$ (6)
Although the illumination part can be obtained via Eqs. 5 and 6, we further incorporate a continuity constraint to align with real-world properties. For a pixel in the video frame, denoted as $p$, we introduce a spatial continuity loss on the predicted $\hat{L}_{n,t}$ as
$\small\mathcal{L}(L_{t})=\mathbb{E}_{t}[v_{t}^{p}\times[\partial_{x}\hat{L}_{n,t}(p)]^{2}+u_{t}^{p}\times[\partial_{y}\hat{L}_{n,t}(p)]^{2}],$ (7)
where $\partial_{x}$ and $\partial_{y}$ are partial derivatives in horizontal and vertical directions, respectively. $v_{t}^{p}$ and $u_{t}^{q}$ are spatially-varying smoothness weights, calculated as
$\small v_{t}^{p}=(\|\partial_{x}\mathcal{T}_{t}(p)\|^{1.2}+\Delta)^{-1},u_{t}^{p}=(\|\partial_{y}\mathcal{T}_{t}(p)\|^{1.2}+\Delta)^{-1},$ (8)
where $\mathcal{T}_{t}$ represents the logarithmic transformation of $I_{d,t}$, and $\Delta$ is a small constant (set to 0.0001) used to avoid division by zero. The diagram for the decomposition implementation is shown in Fig. 2.
3.3 Dual Network for Consistent Decomposition
Overview and motivation. To enhance the consistency of decomposition, it’s crucial for various frames within a video to mutually share features. By incorporating shared features and simultaneous supervisions, the features used for synthesizing Retinex parameters across different frames can exhibit greater consistency. When the video is enhanced frame by frame without feature propagation, the decomposition results tend to be suboptimal due to the inherently challenging nature of maintaining consistency. This observation is empirically validated in our ablation study. Current LLVE methods primarily leverage multiple-frame inputs to facilitate propagation Wang et al. (2021a); Fu et al. (2023a). However, these strategies come with a notable cost in terms of aligning features, potentially affecting inference efficiency.
In this paper, we introduce an efficient strategy for propagation using a dual approach. This approach entails loading two frames, specifically target frame $I_{d,t_{1}}$ and reference frame $I_{d,t_{2}}$ (where $I_{d,t_{2}}$ can be randomly selected from the temporal neighbors of $I_{d,t_{1}}$ during training, and is set as the closest frame of $I_{d,t_{1}}$ during inference), propagating features at the deepest layer of the network through our cross-frame interaction (as depicted in Fig. 3), and concurrently supervising these dual outputs to guarantee consistency.
Figure 4: Visual samples: derive the Retinex decomposition results of enhanced normal-light outputs from the input low-light video data on different datasets.
Feature propagation via long-range attention and short-range fusion in CFIM. Suppose the feature of $I_{d,t_{1}}$ and $I_{d,t_{2}}$ are denoted as $f_{d,t_{1}}$ and $f_{d,t_{2}}$ which is extracted from the same encoder $E$ in the network $f$. To complete the propagation, we initially employ a long-range cross-frame attention operation in the feature space. A traditional attention operation Vaswani et al. (2017) typically consists of the query vector $q$, the key vector $k$, and the value vector $v$. The attention relationship is established using $A(q,k,v)=softmax(q\times k^{T})\times v$, where $\times$ represents matrix multiplication. To enable cross-frame attention, we process the feature via dual paths, as
$\hat{f}_{d,t_{1}}=A(f_{d,t_{1}},f_{d,t_{1}},f_{d,t_{1}})+A(f_{d,t_{1}},f_{d,t_{2}},f_{d,t_{2}}),$ (9)
$\hat{f}_{d,t_{2}}=A(f_{d,t_{2}},f_{d,t_{2}},f_{d,t_{2}})+A(f_{d,t_{2}},f_{d,t_{1}},f_{d,t_{1}}).$
Moreover, a short-range fusion operation is set as the refinement operation to the propagation at both spatial and channel levels. This can be written as
$\small\bar{f}_{d,t_{1}},\bar{f}_{d,t_{2}}=S(C(\hat{f}_{d,t_{1}}\oplus\hat{f}_{d,t_{2}})),$ (10)
where $\oplus$ represents channel concatenation, $C$ refers to the convolution network, and $S$ signifies channel dimension splitting. The decomposition results for $I_{d,t_{1}}$ and $I_{d,t_{2}}$ are obtained by processing $\bar{f}_{d,t_{1}}$ and $\bar{f}_{d,t_{2}}$ through the decoder. We have confirmed that our dual network with synchronous supervision for the outputs of $I_{d,t_{1}}$ and $I_{d,t_{2}}$ can yield superior decomposition and enhancement results compared to individual enhancement strategies.
3.4 Overall Loss Function
During the training, we find it is better to adopt the dual training strategy, i.e., for each $I_{d,t_{1}}$, we sample its neighboring reference $I_{d,t_{2}}$ and constrain their outputs simultaneously. Thus, the loss function can be written as
$\small\mathcal{L}_{v}=\mathbb{E}_{t_{1},t_{2}}[\mathcal{L}_{R,t_{1}}+\mathcal{L}_{R,t_{2}}+\lambda_{1}(\mathcal{L}(L_{t_{1}})+\mathcal{L}(L_{t_{2}}))+\lambda_{2}\mathcal{L}(R_{t_{1},t_{2}})],$ (11)
where $\lambda_{1}$ and $\lambda_{2}$ are the loss weights, and each loss term is defined in Eqs. 5, 6 and 7. Moreover, we observe that the performance of our framework is robust to the choice of the reference frame.
4 Experiments
4.1 Datasets
Our evaluation is conducted on four publicly available datasets, which encompass a wide range of real-world videos with diverse motion patterns and degradations. (1) SMID Chen et al. (2019): This dataset consists of static videos with ground truths obtained through long exposure. We followed the procedure provided by SMID, converting the videos from the RAW domain to the sRGB domain using rawpy’s default ISP. (2) SDSD Wang et al. (2021a): SDSD is a dynamic video dataset collected through electromechanical equipment, including both indoor and outdoor subsets. (3) DID Fu et al. (2023a): DID is another dynamic video dataset characterized by pronounced camera motion, strict spatial alignment, and diverse scene content. In comparison to SDSD, DID offers more data and is a multi-illuminance, multi-camera LLVE dataset. We follow DID’s division into a training set, a test set, and a validation set, maintaining a ratio of 3:1:1. (4) DAVIS Pont-Tuset et al. (2017): DAVIS is a dataset that includes both scene dynamics and camera motions. To simulate low-light degradation in DAVIS, we followed the methodology outlined in Zhang et al. (2021), further incorporating noise degradation.
4.2 Implementation Details
We conducted experiments on all datasets using the same network structure. All modules were trained end-to-end, with the learning rate initialized at $4e^{-4}$ for all layers, adapted by a cosine learning scheduler. The batch size used was 4. Correspondences were computed using the SOTA method DKM Edstedt et al. (2023). Additionally, no data augmentation was applied during training. All experiments were executed on a single NVIDIA A100 GPU.
Table 1: Quantitative comparison on SDSD, SMID, and DID datasets. Our method performs the best consistently.
SDSD-indoor SDSD-outdoor SMID DID
Methods PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM
SNR Xu et al. (2022) 27.30 0.84 23.23 0.82 28.49 0.81 24.85 0.90
SMG Xu et al. (2023b) 27.82 0.86 25.17 0.81 28.03 0.78 25.14 0.85
PairLLE Fu et al. (2023b) 23.48 0.71 20.04 0.65 22.70 0.63 22.56 0.82
RetinexFormer Cai et al. (2023) 26.56 0.79 22.80 0.77 29.15 0.82 25.40 0.89
MBLLEN Lv et al. (2018) 22.17 0.66 21.41 0.63 22.67 0.68 24.22 0.86
SMID Chen et al. (2019) 24.84 0.72 23.30 0.67 24.78 0.72 22.28 0.84
SMOID Jiang and Zheng (2019) 24.63 0.70 22.25 0.68 23.64 0.71 22.13 0.85
SDSDNet Wang et al. (2021a) 26.81 0.75 23.08 0.71 26.03 0.75 22.52 0.81
DP3DF Xu et al. (2023a) 27.63 0.74 23.85 0.73 27.19 0.76 22.39 0.88
StableLLVE Zhang et al. (2021) 25.32 0.70 22.47 0.65 24.37 0.69 21.64 0.80
LLVE-SEG Liu et al. (2023) 26.19 0.77 24.09 0.72 25.31 0.74 23.85 0.82
LLVE-CFA Chhirolya et al. (2022) 24.28 0.81 22.64 0.76 24.72 0.70 22.53 0.87
Ours 28.93 0.88 26.32 0.82 29.60 0.82 30.10 0.93
Table 2: Quantitative comparison on the DAVIS dataset.
Methods SNR SMG PairLLE RetinexFormer MBLLEN SMID
PSNR 20.69 20.18 19.57 21.79 18.63 20.51
SSIM 0.710 0.672 0.667 0.723 0.619 0.686
Methods SMOID SDSDNet DP3DF StableLLVE LLVE-SEG Ours
PSNR 20.66 21.22 22.04 21.48 21.45 23.38
SSIM 0.697 0.716 0.749 0.732 0.715 0.782
4.3 Comparison and Baselines
We conducted a comprehensive comparison with SOTA LLVE methods, including MBLLEN Lv et al. (2018), SMID Chen et al. (2019), SMOID Jiang and Zheng (2019), SDSDNet Wang et al. (2021a), DP3DF Xu et al. (2023a), StableLLVE Zhang et al. (2021), LLVE-SEG Liu et al. (2023), and LLVE-CFA Chhirolya et al. (2022) We also compared our method with SOTA LLIE methods for individual frames, including SNR Xu et al. (2022), SMG Xu et al. (2023b), PairLLE Fu et al. (2023b), RetinexFormer Cai et al. (2023). All these methods were trained on each dataset using their respective released code and hyper-parameters (e.g., the training epoch that is set to guarantee the convergence). Note that all these baselines are trained on our unified data split for a fair comparison. The splits slightly differ from those in baselines’ original papers, so the scores of the baselines may vary from the original papers.
Quantitative result. In Table 1, we present the comparative results with the selected baseline methods across the SDSD, SMID, and DID datasets. The table reveals that our approach consistently outperforms all other methods, as indicated by the highest PSNR and SSIM scores on all datasets. Notably, our scores exhibit a substantial lead over all others, particularly on DID, which is a large-scale dynamic video dataset. This superiority underscores the robust capability of our method in enhancing real-world videos.
Furthermore, Table 2 provides a summary of the comparative results on DAVIS. In comparison to the degradation synthesis strategy as presented in Zhang et al. (2021), we have extended our approach to include the degradation of a low-light noise term. As a result, our synthesized data encompasses dynamic scenes with pronounced invisibility and perturbations, posing a considerable challenge. As illustrated in Table 2, our method consistently yields the highest PSNR and SSIM values, reaffirming the effectiveness of our approach.
Qualitative result. Besides the quantitative comparison, we present visual comparisons with the selected baselines. Fig. 5 showcases the visual comparisons on SDSD and SMID, while Fig. 6 displays visual cases from DID and DAVIS. In general, the results enhanced by our approach exhibit a more natural appearance, including accurate color, well-balanced brightness, enhanced contrast, and precise details without noise. Furthermore, our framework’s results show fewer artifacts in regions with complex textures, and they appear cleaner and sharper compared to the results produced by other methods. This distinction is particularly notable, as most baselines perform well in simpler areas.
Input
RetinexFormer
SDSDNet
DP3DF
Ours
GT
Input
RetinexFormer
SDSDNet
DP3DF
Ours
GT
Input
RetinexFormer
SDSDNet
DP3DF
Ours
GT
Figure 5: Visual comparisons on SDSD-indoor (top), SDSD-outdoor (middle), and SMID (bottom). The results of our proposed framework, i.e., “Ours”, demonstrate better visual perception with less noise, clearer visibility, and more enhanced details.
Input
RetinexFormer
SDSDNet
DP3DF
Ours
GT
Input
RetinexFormer
SDSDNet
DP3DF
Ours
GT
Figure 6: Comparisons on DID (top) and DAVIS (bottom). “Ours” has less noise, better visibility and details.
Figure 7: The above pie charts summarize the results of our user study, and ours is preferred by participants.
4.4 Ablation Study
We assess the critical components of our framework through five ablation cases. (1) “w/o LR” that removes the constraint on learning the reflectance part. (2) “w/o LL” where the constraint on learning the illumination part is removed. (3) “w/o C.F.”: we eliminate the cross-frame attention and fusion in the network, resulting in each frame being enhanced individually. (4) “with M.I.”: This refers to using multiple neighboring frames as input and employing the temporal alignment of deformable convolution. (5) “w/o Dual”: the dual learning strategy is removed, meaning that the loss is applied to only one frame in each iteration.
The results are summarized in Table 3. By comparing “w/o LR” with the full setting (“Full”), we can clearly demonstrate the effectiveness of our proposed constraint for the reflectance map. Similarly, the significant advantage of “Full” over “w/o LL” highlights the significance of the proposed illumination continuity constraint. Furthermore, we can validate the importance of mutually propagating features among different frames to achieve spatial-temporal consistent decomposition by comparing “w/o C.F.” with “Full”. Our proposed propagation strategy via the dual network structure proves to be more effective than the common method of using multi-frame inputs for propagation in videos, as seen in the comparison between “w/o M.I.” and “Full”. The dual learning setting, where we simultaneously supervise the two outputs of the dual network, also proves to be valuable, as evidenced by the comparison between “w/o Dual” and “Full”.
Table 3: Ablation study on SDSD, SMID, and DID.
SDSD-indoor SDSD-outdoor SMID DID
Methods PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM
w/o LR 25.78 0.77 23.49 0.75 26.37 0.76 27.54 0.86
w/o LL 26.89 0.80 24.19 0.78 27.74 0.79 27.72 0.89
w/o C.F. 25.18 0.82 24.22 0.77 26.75 0.77 25.60 0.85
with M.I. 26.21 0.84 25.36 0.80 28.42 0.79 28.33 0.88
w/o Dual 27.53 0.84 24.83 0.79 27.43 0.78 26.91 0.86
Full 28.93 0.88 26.32 0.82 29.60 0.82 30.10 0.93
4.5 User Study
To prove the effectiveness of our proposed framework in terms of human subjective evaluation, we conduct a large-scale user study with 100 participants with varying ages and education backgrounds, and balanced sex distribution. We aim to compare the perceptual quality of our method against SOTA baselines. To conduct the user study, we randomly collect 30 videos in the wild with the capture device of Sony A7R3 and iPhone13 (all models in the user study are trained with DID, since it is a large-scale dataset with multiple scenes and devices). Following most of the existing low-light enhancement works Xu et al. (2023b); Wang et al. (2023a), we employ the AB-test for the user study. For a video, the result produced by our method is called “Video A” whereas the result from another baseline is “Video B”. During the evaluation, each participant will compare the results of our method against other baselines in random order. For the comparison with one baseline, 5 videos will be randomly selected for visual perception. Each participant would simultaneously watch videos A and B. To avoid the bias caused by the position, we will randomize the left-right presentation order when showing videos A and B in each question. The participants can choose among three options: “Video A is better”, “Video B is better”, and “I think they are of the same quality”. They should make the decision according to the natural brightness, contrast, and color of each frame; the rich details, fewer artifacts, and the temporal consistency in the video. For each participant, the number of tasks is 10 methods $\times$ 5 videos $=50$, and it took around 50 minutes on average for each participant to complete the user study.
Fig. 7 summarizes the user study’s results, and we can see that ours gets more selections from participants over all the baselines. This demonstrates that our method’s results are more preferred by the human subjective perception.
Table 4: Quantitative comparison on 3D low-light datasets with normal-light pairs, proposed in Wang et al. (2023a).
Methods PairLLE SNR RetinexFormer SCI+NeRF Ours
PSNR 18.35 18.87 19.49 13.08 21.09
SSIM 0.604 0.650 0.702 0.505 0.763
Methods PairLLE+NeRF SNR+NeRF RetinexFormer+NeRF LLNeRF Ours+NeRF
PSNR 18.72 19.24 19.98 20.50 22.35
SSIM 0.681 0.693 0.736 0.758 0.782
4.6 Enhancement for NeRF
Recent works have introduced the implementation of 3D low-light enhancement Wang et al. (2023a), with a primary focus on disentangling view-independent intrinsic color and view-varying illuminations to achieve 3D-consistent enhancement. On the other hand, our framework’s enhanced results are particularly well-suited for NeRF reconstruction, as we have implemented a decomposition that maintains consistency across different spatial locations within a 3D scene.
To assess the impact of using our enhanced results for NeRF, we conducted experiments on a multi-view low-light dataset previously introduced in Wang et al. (2023a). We employed the network trained on DID, as it benefits from a large-scale training dataset, resulting in solid generalization capabilities. Furthermore, DID inherently contains multi-view images. Specifically, for each view, we selected its closest view as the reference view for our input data.
In Table 4, we present the PSNR and SSIM results on the testing set of Wang et al. (2023a), where ground truth data is available. It’s important to note that none of the networks were fine-tuned on the testing set. The superiority of our framework is evident, as our network can achieve 3D-consistent enhancement that aligns with the fundamental principles of NeRF.
5 Conclusion
We introduce a novel LLVE framework that utilizes spatial-temporal consistent illumination and reflection decomposition and dual networks for mutual feature propagation. It predicts the decomposition of normal-light outputs without external priors, thanks to cross-frame correspondences and practical illumination continuity. Extensive experiments on various datasets showcase the framework’s superiority over SOTA approaches and the impact of our designed components.

### v2

@@
-Low-Light Video Enhancement via Spatial-Temporal Consistent Illumination and Reflection Decomposition
+Low-Light Video Enhancement via Spatial-Temporal Consistent Decomposition
 Abstract
-Low-Light Video Enhancement (LLVE) seeks to restore dynamic and static scenes plagued by severe invisibility and noise.
-One critical aspect is formulating a consistency constraint specifically for temporal-spatial illumination and appearance enhanced versions, a dimension overlooked in existing methods.
-In this paper, we present an innovative video Retinex-based decomposition strategy that operates without the need for explicit supervision to delineate illumination and reflectance components.
-We leverage dynamic cross-frame correspondences for intrinsic appearance and enforce a scene-level continuity constraint on the illumination field to yield satisfactory consistent decomposition results.
-To further ensure consistent decomposition, we introduce a dual-structure enhancement network featuring a novel cross-frame interaction mechanism.
+Low-Light Video Enhancement (LLVE) seeks to restore dynamic or static scenes plagued by severe invisibility and noise.
+In this paper, we present an innovative video decomposition strategy that incorporates view-independent and view-dependent components to enhance the performance of LLVE.
+We leverage dynamic cross-frame correspondences for the view-independent term (which primarily captures intrinsic appearance) and impose a scene-level continuity constraint on the view-dependent term (which mainly describes the shading condition) to achieve consistent and satisfactory decomposition results.
+To further ensure consistent decomposition, we introduce a dual-structure enhancement network featuring a cross-frame interaction mechanism.
+By supervising different frames simultaneously, this network encourages them to exhibit matching decomposition features.
 This mechanism can seamlessly integrate with encoder-decoder single-frame networks, incurring minimal additional parameter costs.
-By supervising different frames simultaneously, this network encourages them to exhibit matching decomposition features, thus achieving the desired temporal propagation.
 Extensive experiments are conducted on widely recognized LLVE benchmarks, covering diverse scenarios.
-Our framework consistently outperforms existing methods, establishing a new state-of-the-art (SOTA) performance.
+Our framework consistently outperforms existing methods, establishing a new SOTA performance.
 1 Introduction
-Low-light enhancement aims to enhance underexposed images and videos captured in low-light conditions Xu et al.
-(2022); Wang et al.
-(2021a), improving their visual quality while reducing noise.
-This is a longstanding and important topic in computer vision, with wide-ranging applications, such as improving portrait photography on mobile devices Ignatov et al.
-(2017); Hasinoff et al.
-(2016) and supporting various downstream tasks, including nighttime face recognition Ma et al.
-(2022); Wang et al.
-(2022a) and vehicle detection Fu et al.
-(2023a); Wu et al.
-(2022b).
-Figure 1: Our method consistently achieves SOTA performance on different LLVE datasets with the same network architecture.
-Numerous methods have been proposed to enhance underexposed images using supervised learning Wang et al.
-(2021b); Xu et al.
-(2022, 2023b); Cai et al.
-(2023); Wang et al.
-(2022b), yielding impressive results.
-However, these methods still struggle to provide satisfactory performance when applied to videos, a popular media format.
+Low-light enhancement aims to enhance underexposed images and videos captured in low-light conditions (Xu et al., 2022; Wang et al., 2021), improving their visual quality while reducing noise.
+This technique can be applied in wide-ranging applications, such as portrait photography on mobile devices (Ignatov et al., 2017; Hasinoff et al., 2016), nighttime face recognition (Ma et al., 2022; Wang et al., 2022) and vehicle detection (Fu et al., 2023a; Wu et al., 2022).
 The key challenge when enhancing videos is the need for consistent enhancement results across corresponding locations in different frames, which vary both spatially and temporally.
-Current methods rely on paired training data, where the ground truth ensures this consistency.
-Unfortunately, there is a scarcity of real-world, high-quality, spatially aligned video pairs for dynamic scenes, resulting in limited generalization capabilities for unseen videos with varying conditions.
-One solution is to incorporate parameterized enhancement models with physical significance into the enhancement process, reducing the reliance on training data, e.g., Retinex theory Land (1977).
-According to this theory, an observed image $I$ can be decomposed into its illumination component $L$ and reflectance component $R$ using the equation $I=L\otimes R$, where $\otimes$ denotes element-wise multiplication.
-Several approaches Zhang et al.
-(2019); Wu et al.
-(2022a); Fu et al.
-(2023b) have been developed to predict both $R$ and $L$ under normal illumination conditions for a low-light input image.
-However, a suitable strategy for obtaining the Retinex decomposition in normal-light conditions for low-light input videos is currently lacking.
-Ensuring both temporal and spatial consistency in decomposition is crucial, yet previous strategies were primarily designed for individual images, thus posing a challenge for video processing.
-In this paper, we introduce an innovative approach to attain a consistent Retinex decomposition for normal-light video outputs, ensuring consistency in both spatial and temporal dimensions.
+Moreover, there is a scarcity of real-world, high-quality, spatially aligned video pairs for dynamic scenes, resulting in limited generalization capabilities for unseen videos with varying conditions.
+Figure 1: Our proposed LLVE method consistently achieves SOTA performance on different LLVE datasets involving various scenes with the same network architecture.
+One solution is to incorporate parameterized enhancement models with physical significance into the enhancement process, reducing the reliance on training data.
+According to the linear Lambertian model for intrinsic image decomposition (Yuille, 2012; Narihira et al., 2015), an observed image $\boldsymbol{I}$ can be expressed as the element-wise product ($\otimes$) of its albedo component and shading component.
+The albedo part is normally view-independent, capturing the intrinsic appearance, whereas the shading component is commonly view-dependent, influenced by lighting conditions and the surface normal’s direction.
+Inspired by the Lambertian model, we propose that a frame in the LLVE task can similarly be decomposed as $\boldsymbol{I}=\boldsymbol{L}\otimes\boldsymbol{R}$, where $\boldsymbol{L}$ represents the view-dependent term, and $\boldsymbol{R}$ corresponds to the view-independent component.
+In this paper, we introduce an innovative approach to attain our decomposition for normal-light video outputs, ensuring consistency in both spatial and temporal dimensions.
 We leverage cross-frame correspondences and real-world physical continuity constraints to achieve this decomposition without the need for explicit supervision.
 Furthermore, to enhance the consistency of the decomposition, we’ve developed an efficient dual-structure enhancement network featuring a novel cross-frame interaction mechanism.
-Differing from some previous methods Wang et al.
-(2019a, 2021a), which directly assume the reflection term as the enhanced version, our approach follows Cai et al.
-(2017); Wei et al.
-(2018); Wu et al.
-(2022a); Fu et al.
-(2023b), i.e., aims to recover both the reflection and illumination components of normal-light images from the provided low-light inputs.
-Previous methods have verified that satisfying image-level reflection and illumination decomposition can be achieved with suitable reflection and illumination priors via a data-driven manner.
-In this paper, we aim to design appropriate priors for video data to implement the decomposition.
-We recognize that the reflectance component $R$ should represent the physical properties of objects, which should remain consistent across various observation perspectives Wei et al.
-(2018).
-To achieve this, we first calculate corresponding locations in videos and utilize these computed spatial-temporal correspondences, along with uncertainty values, to link the reflection terms across different frames.
-Conversely, the illumination terms are subject to a spatial smoothness loss to model real-world physical constraints Wei et al.
-(2018).
-This approach allows us to establish a consistency constraint specifically for temporal-spatial illumination and appearance in the enhanced videos, without requiring additional supervision information.
-Additionally, the architecture of the enhancement networks plays a crucial role in shaping the decomposition results.
-It is important that different frames share features and receive simultaneous supervision, promoting the alignment of feature representations to synthesize Retinex parameters.
-In our paper, we introduce a novel dual network structure capable of achieving such feature propagation.
-Two frames are used as inputs, and their features are shared within the designed Cross-Frame Interaction Module (CFIM).
+Our approach aims to predict both the $\boldsymbol{R}$ and $\boldsymbol{L}$ of normal-light images from the provided low-light inputs.
+It is important to note that the decomposition of view-independent and view-dependent terms is not unique, and our goal is to implement a suitable decomposition that enhances the performance of LLVE.
+In this paper, we aim to design appropriate priors directly from the video data to implement the decomposition.
+We recognize that $\boldsymbol{R}$ should represent the physical properties of objects, which should remain consistent across various observation perspectives.
+To achieve this, we first calculate corresponding locations in videos and utilize these computed spatial-temporal correspondences, along with uncertainty values, to link the predictions of $\boldsymbol{R}$ across different frames.
+Conversely, the terms of $\boldsymbol{L}$ are subject to a spatial smoothness loss to model real-world physical constraints (since the lighting field is continuous in practice).
+This approach allows us to establish a temporal-spatial consistency constraint in the enhanced videos, without requiring additional supervision.
+Note that we cannot guarantee the decomposition results are strictly view-independent or view-dependent, as this is not a 3D method.
+Our goal is to apply appropriate constraints to approximate desired decomposition properties, while focusing on optimizing the final LLVE performance with our approach.
+Additionally, we introduce a dual-network structure to enforce the consistency of feature representations across frames, facilitating the consistent synthesis of decomposition in turn.
+Two frames are used as inputs, and their features are propagated within the designed Cross-Frame Interaction Module (CFIM).
+CFIM differs from other similar architectures in the restoration task (Lv et al., 2023) in terms of the fusion manner.
 CFIM facilitates interaction through a combination of long-range cross-frame attention computation and short-range channel-spatial fusion, ensuring the accurate and comprehensive merging of features from both global and local perspectives.
-CFIM can be seamlessly integrated into any encoder-decoder single-image enhancement structure by inserting it at the deepest hidden layer.
-In this paper, we illustrate this with the simplest U-Net Ronneberger et al.
-(2015) as an example, and it’s worth noting that CFIM comprises only 0.4M parameters.
-We performed experiments on various widely recognized video enhancement datasets Wang et al.
-(2021a); Chen et al.
-(2019); Fu et al.
-(2023a); Zhang et al.
-(2021).
+Moreover, CFIM selectively incorporates knowledge from randomly chosen neighboring frames into the backbone (i.e., the part without feature propagation) during training.
+In this way, CFIM help the backbone in learning how to adaptively utilize knowledge from various cross-frame features, resulting in a more robust backbone.
+We performed experiments on various widely recognized video enhancement datasets.
 Our results, both quantitative and qualitative, consistently demonstrate the effectiveness and state-of-the-art (SOTA) performance of our framework, as shown in Fig.
 1.
-Furthermore, we conducted a large-scale user study involving 100 participants, which clearly showcased the superiority of our enhanced results in terms of human subjective perceptions.
+Furthermore, we conducted a large-scale user study involving 100 participants, which showcased the superiority of our results in terms of human subjective perceptions.
 In summary, our contributions are three-fold.
-  • We introduce a novel canonical form for LLVE that predicts spatial-temporal consistent Retinex decomposition for normal-light outputs.
-This innovative decomposition strategy leverages spatial-temporal correspondences and continuity.
-  • We design a new LLVE network that facilitates interaction among the features of different frames, leading to a consistent decomposition.
+  • We introduce a novel canonical form for LLVE that predicts spatial-temporal consistent decomposition with view-independent and view-dependent terms for normal-light outputs.
+This decomposition strategy leverages spatial-temporal correspondences and continuity.
+  • We design a new LLVE network that facilitates interaction among the features of different frames and fits our consistent decomposition strategy.
   • We conduct extensive experiments on public datasets, illustrating the effectiveness of our proposed framework.
@@
 To enhance the quality of a low-light video, image enhancement methods can be applied on a frame-by-frame basis.
-In recent years, learning-based Low-Light Image Enhancement (LLIE) techniques Yan et al.
-(2014, 2016); Lore et al.
-(2017); Cai et al.
-(2018); Zamir et al.
-(2020); Xu et al.
-(2020); Zeng et al.
-(2020); Kim et al.
-(2021); Zhao et al.
-(2021); Zheng et al.
-(2021); Wang et al.
-(2021b); Liu et al.
-(2021); Yang et al.
-(2021b); Jiang et al.
-(2021); Yang et al.
-(2021a) have made significant advancements, with a primary focus on supervised approaches thanks to the abundance of image pairs available for training.
-Xu et.al.
-Xu et al.
-(2022) introduced spatial-varying operations, considering Signal-to-Noise Ratio (SNR) as a prior factor.
-However, when applying image enhancement algorithms to individual frames, the issue of flickering often arises.
+In recent years, learning-based Low-Light Image Enhancement (LLIE) techniques (Jiang et al., 2021; Yang et al., 2021) have made significant advancements, with a primary focus on supervised approaches.
 Low-Light Video Enhancement.
 In addition to the need for LLIE, there is a growing demand for video enhancement, considering the widespread use of videos as a popular data format on the internet and in photographic equipment.
-Various approaches have been proposed in this context Chen et al.
-(2019); Triantafyllidou et al.
-(2020); Ye et al.
-(2023); Lv et al.
-(2023); Xu et al.
-(2023a); Fu et al.
-(2023a).
-Danai et.al.
-Triantafyllidou et al.
-(2020) introduced a data synthesis mechanism to generate dynamic video pairs from SID Chen et al.
-(2019).
-Wang et.al.
-Wang et al.
-(2021a) proposed a multi-branch network that simultaneously estimates noise and illumination, suitable for videos with severe noise conditions.
+Various approaches have been proposed in this context (Chen et al., 2019; Triantafyllidou et al., 2020; Ye et al., 2023; Lv et al., 2023; Xu et al., 2023a; Fu et al., 2023a).
 Liu et.al.
-Liu et al.
-(2023) and Liang et.al.
-Liang et al.
-(2023) used prior event information to learn enhancement mapping for brightening videos.
+(Liu et al., 2023) and Liang et.al.
+(Liang et al., 2023) used prior event information to learn enhancement mapping for brightening videos.
 Xu et.al.
-Xu et al.
-(2023a) designed a parametric 3D filter tailored for enhancing and sharpening low-light videos.
+(Xu et al., 2023a) designed a parametric 3D filter tailored for enhancing and sharpening low-light videos.
 Recently, Fu et.al.
-Fu et al.
-(2023a) introduced a video enhancement method called LAN, which iteratively refines illumination and adaptively adjusts it.
+(Fu et al., 2023a) introduced a video enhancement method called LAN, which iteratively refines illumination and adaptively adjusts it.
 However, it’s important to note that LAN lacks an explicit constraint for maintaining consistent reflectance and illumination decomposition.
-Several datasets for video enhancement, encompassing both static Chen et al.
-(2019); Jiang and Zheng (2019); Wang et al.
-(2019b); Triantafyllidou et al.
-(2020) and dynamic motions Wang et al.
-(2021a); Fu et al.
-(2023a), have been introduced.
+Several datasets for video enhancement, encompassing both static (Chen et al., 2019; Jiang and Zheng, 2019; Wang et al., 2019; Triantafyllidou et al., 2020) and dynamic motions (Wang et al., 2021; Fu et al., 2023a), have been introduced.
 In this paper, we introduce a novel LLVE method, designed to enhance effects on these datasets.
-Our decomposition method explicitly and consistently models reflectance and illumination for all frames, with the guidance of the videos themselves.
-3 Method
-We commence by revisiting the formulation of the Retinex theory in Sec.
-3.1.
-Following that, we provide a detailed exposition of our decomposition strategy specifically tailored for video data in Sec.
-3.2.
-Subsequently, in Sec.
-3.3, we delve into the description of our cross-frame interaction.
-Finally, the training loss for our framework will be introduced in detail.
-Figure 2: Our framework provides a comprehensive solution that includes explicit and consistent modeling for Retinex decomposition of enhanced normal-light outputs across different frames.
-To achieve this, we enforce consistent features in the reflectance terms across different frames by leveraging computed correspondences in the temporal dimension of videos ($\mathcal{L}(R_{t})$).
-Simultaneously, we ensure that the illumination terms exhibit a spatially continuous distribution ($\mathcal{L}(L_{t})$), aligning with real-world scenarios.
-Furthermore, our network incorporates cross-frame interaction and simultaneous supervision of different frames within a video, encouraging consistent Retinex features among these frames.
+Our decomposition method explicitly and consistently models view-dependent and -independent for all frames.
+Figure 2: Our framework offers a comprehensive solution that explicitly and consistently models the view-independent and view-dependent decomposition of enhanced normal-light outputs across different frames.
+To achieve this, we enforce consistent features in the view-independent terms across different frames by leveraging computed correspondences in the temporal dimension of videos ($\mathcal{O}(\boldsymbol{R}_{t})$).
+Simultaneously, we ensure that the view-dependent terms exhibit a spatially continuous distribution ($\mathcal{O}(\boldsymbol{L}_{t})$), aligning with real-world scenarios.
+Furthermore, our network incorporates cross-frame interaction and simultaneous supervision of different frames within a video, encouraging consistent features for these frames derived from one video.
 For a more detailed visual representation, please refer to Fig.
 3.
-3.1 Retinex Model
-Retinex Theory.
-When presented with an image $I$ (in this paper, we denote the low-light image as $I_{d}$ and the normal-light image as $I_{n}$, with $d$ and $n$ serving as subscript abbreviations), we assume its decomposition into illumination $L$ and reflectance $R$ follows the Retinex theory Land and McCann (1971); Fu et al.
-(2023b), as
-$I=L\otimes R,$ (1)
-where $\otimes$ denotes the element-wise multiplication, $L$ and $R$ can be formulated for different channels Wang et al.
-(2019a); Cai et al.
-(2023), i.e., the channel number is 3 for the sRGB domain.
-In this context, $L$ describes the light intensity of objects, which is expected to exhibit piece-wise continuity.
-On the other hand, $R$ represents the physical properties of the objects, encompassing textures and details observed in the data $I$.
-If we obtain $L$ and $R$ in the normal-light conditions for a given input low-light image, the target normal-light image can be acquired accordingly.
-Numerous methods have been proposed to address this decomposition problem Zhang et al.
-(2019); Wu et al.
-(2022a); Fu et al.
-(2023b), and their common objective can be summarized as
-$\mathcal{L}=\|I-L\otimes R\|+\mathcal{L}(L)+\mathcal{L}(R),$ (2)
-where $\mathcal{L}(L)$ and $\mathcal{L}(R)$ denote the constraints for illumination $L$ and reflectance $R$, respectively.
-Shortcoming of current Retinex models for videos.
-Several prior approaches have been devised to attain this desired decomposition by imposing different constraints on the two components, such as structural and smoothness constraints Li et al.
-(2018); Wei et al.
-(2018).
-Nevertheless, there is a scarcity of satisfactory Retinex decomposition prediction tailored for sequential video data.
-The issue concerning the decomposition of video data can be summarized as
-$\mathcal{L}_{v}=\mathbb{E}_{t=1:T}[\mathcal{L}_{R,t}+\mathcal{L}(L_{t})+\mathcal{L}(R_{t})],\quad\mathcal{L}_{R,t}=\|I_{t}-L_{t}\otimes R_{t}\|,$ (3)
-where $\mathbb{E}$ is the average operation, $T$ is the number of frame.
-Compared with Eq.
-2, we need to simultaneously guarantee the decomposition quality and the consistency among different frames’ results.
-Video data guide Retinex models by themselves.
-We have discovered that video data itself can play a crucial role in facilitating the achievement of a satisfactory Retinex decomposition prediction for normal-light outputs, thus obviating the need for external priors typically used in low-light enhancement.
-By establishing correspondences among different frames, we can impose specific constraints: the reflectance of each frame, denoted as $R_{t}$, should faithfully represent the intrinsic texture of objects within the target scene, remaining consistent regardless of changes in viewpoint.
-Likewise, the illumination of each frame, denoted as $L_{t}$, should exhibit the desired continuity consistent with real-world physical properties.
-For further elaboration, please refer to Sec.
-3.2.
+3 Method
+3.1 Decomposition Model
+Motivation.
+According to the linear Lambertian model (Narihira et al., 2015), an observed image $\boldsymbol{I}$ can be formulated as the element-wise product of its albedo and shading component.
+Here, the albedo part is view-independent, capturing the intrinsic appearance, while the shading component is view-dependent.
+Inspired by this, when presented with an image $\boldsymbol{I}$ (in this paper, we denote the low-light image as $\boldsymbol{I}_{d}$ and the normal-light image as $\boldsymbol{I}_{n}$, with $d$ and $n$ serving as subscript abbreviations), we assume its decomposition into view-dependent part $\boldsymbol{L}$ and view-independent part $\boldsymbol{R}$, as
+$\boldsymbol{I}=\boldsymbol{L}\otimes\boldsymbol{R},$ (1)
+where $\otimes$ denotes the element-wise multiplication, $\boldsymbol{L}$ and $\boldsymbol{R}$ can be formulated for different channels, i.e., the channel number is 3 for the sRGB domain.
+In this context, $\boldsymbol{L}$ mainly describes the light intensity of objects, which is expected to exhibit piece-wise continuity.
+On the other hand, $\boldsymbol{R}$ primarily represents the physical properties of the objects, encompassing textures and details observed in the data $\boldsymbol{I}$.
+If we obtain $\boldsymbol{L}$ and $\boldsymbol{R}$ in the normal-light conditions for a given input low-light image, the target normal-light image can be acquired accordingly.
+The common objective can be summarized as
+$\mathcal{O}=\|\boldsymbol{I}-\boldsymbol{L}\otimes\boldsymbol{R}\|+\mathcal{O}(\boldsymbol{L})+\mathcal{O}(\boldsymbol{R}),$ (2)
+where $\mathcal{O}(\boldsymbol{L})$ and $\mathcal{O}(\boldsymbol{R})$ denote the constraints for $\boldsymbol{L}$ and $\boldsymbol{R}$ (we will introduce our designed new constraints in Sec.
+3.2).
+When applying this model to sequential video data, the objective related to the decomposition of video data can be summarized as
+$\mathcal{O}_{v}=\mathbb{E}_{t=1:T}[\mathcal{O}_{rec,t}+\mathcal{O}(\boldsymbol{L}_{t})+\mathcal{O}(\boldsymbol{R}_{t})],$ (3)
+$\mathcal{O}_{rec,t}=\|\boldsymbol{I}_{t}-\boldsymbol{L}_{t}\otimes\boldsymbol{R}_{t}\|,$
+where $\mathbb{E}$ is the average operation, $T$ is the number of frame, and $\boldsymbol{L}_{t}$/$\boldsymbol{R}_{t}$ is the decomposition output at the time index of $t$.
+Compared with Equation 2, we need to simultaneously guarantee the decomposition quality and the consistency among different frames’ results.
+Video Data Guide the Decomposition by Themselves.
+We have discovered that video data itself can play a crucial role in facilitating the achievement of our decomposition predictions for normal-light outputs, thus obviating the need for external priors typically used in low-light enhancement.
+By establishing correspondences among different frames, we can impose specific constraints: the $\boldsymbol{R}_{t}$ of each frame should faithfully represent the intrinsic texture of objects within the target scene, remaining consistent regardless of changes in viewpoint.
+Likewise, the $\boldsymbol{L}_{t}$ of each frame should exhibit the desired continuity consistent with real-world physical properties.
+For further elaboration, please refer to Section 3.2.
 Problem Formulation.
 We adopt the supervised setting.
-Given a clip of low-light data $I_{d,t},t\in[1,T]$, there is a paired normal-light data $I_{n,t},t\in[1,T]$.
-We aim to directly obtain the decomposition of $I_{n,t}$ from $I_{d,t}$ using network $f$, as
-$\hat{L}_{n,t},\hat{R}_{n,t}=f(I_{d,t}),\hat{I}_{n,t}=\hat{L}_{n,t}\otimes\hat{R}_{n,t},$ (4)
-where $\hat{L}_{n,t}$ and $\hat{R}_{n,t}$ are the estimated targets (several examples are displayed in Fig. 4), and $\hat{I}_{n,t}$ is the predicted enhancement result.
-Figure 3: The lightweight cross-frame interaction mechanism (CFIM) to propagate different frames’ features, along with the cross-frame attention and spatial-channel fusion.
-Cross-frame interaction can be employed in the deep feature space of arbitrary single-image encoder-decoder frameworks, and we choose U-Net here.
+Given a clip of low-light data $\boldsymbol{I}_{d,t},t\in[1,T]$, there is a paired normal-light data $\boldsymbol{I}_{n,t},t\in[1,T]$.
+We aim to directly obtain the decomposition of $\boldsymbol{I}_{n,t}$ from $\boldsymbol{I}_{d,t}$ using network $f$, as
+$\hat{\boldsymbol{L}}_{n,t},\hat{\boldsymbol{R}}_{n,t}=f(\boldsymbol{I}_{d,t}),\hat{\boldsymbol{I}}_{n,t}=\hat{\boldsymbol{L}}_{n,t}\otimes\hat{\boldsymbol{R}}_{n,t},$ (4)
+where $\hat{\boldsymbol{L}}_{n,t}$ and $\hat{\boldsymbol{R}}_{n,t}$ are the estimated targets, and $\hat{\boldsymbol{I}}_{n,t}$ is the predicted enhancement result.
+Our framework is shown in Fig.
+2 that will be introduced in the following sections.
 3.2 Spatial-Temporal Consistent Decomposition
 Motivation.
-Video can be conceptualized as static/dynamic 3D multi-view data Wang et al.
-(2023b, a).
-Prior research has demonstrated that the multi-view data itself can be harnessed to decompose the view-dependent and view-independent elements, which respectively encapsulate the intrinsic texture and view-altering illuminations Wang et al.
-(2023a).
-Thus, once we establish correspondence relationships among $I_{d,t}$ for all $t$, we can apply the view-independent constraint to derive the reflectance component, denoted as $\hat{R}_{n,t}$, as it represents the intrinsic properties of the target scene.
-Subsequently, obtaining the illumination component, denoted as $\hat{L}_{n,t}$, becomes achievable through the utilization of the inherent image reconstruction loss, in conjunction with adherence to the continuity assumption.
+Video can be conceptualized as static/dynamic 3D multi-view data (Wang et al., 2023b, a).
+Prior research has demonstrated that the multi-view data itself can be harnessed to decompose the view-dependent and view-independent elements, which respectively encapsulate the intrinsic texture and view-altering illuminations (Wang et al., 2023a).
+Thus, once we establish correspondence relationships among $\boldsymbol{I}_{d,t}$ for all $t$, we can apply the view-independent constraint to derive $\hat{\boldsymbol{R}}_{n,t}$, as it represents the intrinsic properties of the target scene.
+Subsequently, obtaining $\hat{\boldsymbol{L}}_{n,t}$ becomes achievable through the utilization of the inherent image reconstruction loss, in conjunction with adherence to the continuity assumption.
 Implementation of Correspondences.
-Obtaining correspondences for $I_{d,t}$, $t\in[1,T]$, can be a challenging task due to the presence of various degradations, including visibility issues and noise.
+Obtaining correspondences for $\boldsymbol{I}_{d,t}$, $t\in[1,T]$, can be a challenging task due to the presence of various degradations, including visibility issues and noise.
 Fortunately, we have access to corresponding normal-light data, which significantly aids in establishing these correspondences, as illustrated in Fig.
-2, where $I_{d,t}$ and $I_{n,t}$ are pixel-wise aligned.
-Given two frames $I_{n,t_{1}}$ and $I_{n,t_{2}}$, correspondences can be determined using the prediction network $F$ Edstedt et al.
-(2023).
+2, where $\boldsymbol{I}_{d,t}$ and $\boldsymbol{I}_{n,t}$ are pixel-wise aligned.
+Given two frames $\boldsymbol{I}_{n,t_{1}}$ and $\boldsymbol{I}_{n,t_{2}}$, correspondences can be determined using the prediction network $F$ (Edstedt et al., 2023).
 We assume the detection of $M$ correspondences, denoted as ${c_{m}=(x_{1,m},y_{1,m},x_{2,m},y_{2,m}),m\in[1,M]}$.
-Each correspondence, represented by $c_{m}$, consists of four coordinates: $(x_{1,m},y_{1,m})$ represents the pixel coordinate in $I_{n,t_{1}}$, while $(x_{2,m},y_{2,m})$ signifies the coordinate in $I_{n,t_{2}}$.
+Each correspondence, represented by $c_{m}$, consists of four coordinates: $(x_{1,m},y_{1,m})$ represents the pixel coordinate in $\boldsymbol{I}_{n,t_{1}}$, while $(x_{2,m},y_{2,m})$ signifies the coordinate in $\boldsymbol{I}_{n,t_{2}}$.
 Furthermore, each correspondence is associated with an uncertainty value $u_{m}$, determined through a data-driven approach.
-Constraints formulation. To satisfy the constraint that the corresponding pixels in the reflectance part are matched, we can formulate $\mathcal{L}(R_{t})$ as
-$\small\mathcal{L}(R_{t_{1},t_{2}})=\mathbb{E}_{m\in[1,M]}u_{m}\|\hat{R}_{n,t_{1}}[x_{1,m},y_{1,m}]-\hat{R}_{n,t_{2}}[x_{2,m},y_{2,m}]\|.$ (5)
-The reconstruction item in Eq. 3 can be denoted as
-$\small\mathcal{L}_{R,t}=\|I_{n,t}-\hat{L}_{n,t}\otimes\hat{R}_{n,t}\|.$ (6)
-Although the illumination part can be obtained via Eqs.
-5 and 6, we further incorporate a continuity constraint to align with real-world properties.
-For a pixel in the video frame, denoted as $p$, we introduce a spatial continuity loss on the predicted $\hat{L}_{n,t}$ as
-$\small\mathcal{L}(L_{t})=\mathbb{E}_{t}[v_{t}^{p}\times[\partial_{x}\hat{L}_{n,t}(p)]^{2}+u_{t}^{p}\times[\partial_{y}\hat{L}_{n,t}(p)]^{2}],$ (7)
+Figure 3: The lightweight cross-frame interaction mechanism (CFIM) propagates different frames’ features, along with cross-frame attention and spatial-channel fusion.
+Cross-frame interaction can be employed in the deep feature space of arbitrary single-image encoder-decoder frameworks, and we choose U-Net here.
+Constraints Formulation. To satisfy the constraint that the corresponding pixels in $\boldsymbol{R}_{t}$ are matched, we have $\mathcal{O}(\boldsymbol{R}_{t})$ as
+$\small\mathcal{O}(\boldsymbol{R}_{t_{1},t_{2}})=\mathop{\mathbb{E}}\limits_{m\in[1,M]}u_{m}\|\hat{\boldsymbol{R}}_{n,t_{1}}[x_{1,m},y_{1,m}]-\hat{\boldsymbol{R}}_{n,t_{2}}[x_{2,m},y_{2,m}]\|.$ (5)
+The reconstruction item in Equation 3 can be denoted as
+$\mathcal{O}_{rec,t}=\|\boldsymbol{I}_{n,t}-\hat{\boldsymbol{L}}_{n,t}\otimes\hat{\boldsymbol{R}}_{n,t}\|.$ (6)
+Although the view-dependent part can be obtained via the reconstruction constraint, We further incorporate a continuity constraint to align with real-world properties.
+For a pixel in the video frame, denoted as $p$, we introduce a spatial continuity objective on the predicted $\hat{\boldsymbol{L}}_{n,t}$ as the following loss function
+$\mathcal{O}(\boldsymbol{L}_{t})=\mathbb{E}_{t}[v_{t}^{p}\times[\partial_{x}\hat{\boldsymbol{L}}_{n,t}(p)]^{2}+u_{t}^{p}\times[\partial_{y}\hat{\boldsymbol{L}}_{n,t}(p)]^{2}],$ (7)
 where $\partial_{x}$ and $\partial_{y}$ are partial derivatives in horizontal and vertical directions, respectively. $v_{t}^{p}$ and $u_{t}^{q}$ are spatially-varying smoothness weights, calculated as
-$\small v_{t}^{p}=(\|\partial_{x}\mathcal{T}_{t}(p)\|^{1.2}+\Delta)^{-1},u_{t}^{p}=(\|\partial_{y}\mathcal{T}_{t}(p)\|^{1.2}+\Delta)^{-1},$ (8)
-where $\mathcal{T}_{t}$ represents the logarithmic transformation of $I_{d,t}$, and $\Delta$ is a small constant (set to 0.0001) used to avoid division by zero.
-The diagram for the decomposition implementation is shown in Fig.
+$\small v_{t}^{p}=(\|\partial_{x}\boldsymbol{U}_{t}(p)\|+\Delta)^{-1},u_{t}^{p}=(\|\partial_{y}\boldsymbol{U}_{t}(p)\|+\Delta)^{-1},$ (8)
+where $\boldsymbol{U}_{t}$ represents the logarithmic transformation of $\boldsymbol{I}_{d,t}$, and $\Delta$ is a small constant (set to 0.0001) used to avoid division by zero.
+The diagram is shown in Fig.
 2.
@@
 To enhance the consistency of decomposition, it’s crucial for various frames within a video to mutually share features.
-By incorporating shared features and simultaneous supervisions, the features used for synthesizing Retinex parameters across different frames can exhibit greater consistency.
+By incorporating shared features and simultaneous supervision, the features used for synthesizing decomposition parameters across different frames can exhibit greater consistency.
 When the video is enhanced frame by frame without feature propagation, the decomposition results tend to be suboptimal due to the inherently challenging nature of maintaining consistency.
 This observation is empirically validated in our ablation study.
-Current LLVE methods primarily leverage multiple-frame inputs to facilitate propagation Wang et al.
-(2021a); Fu et al.
-(2023a).
+Current LLVE methods primarily leverage multiple-frame inputs to facilitate propagation (Wang et al., 2021; Fu et al., 2023a).
 However, these strategies come with a notable cost in terms of aligning features, potentially affecting inference efficiency.
 In this paper, we introduce an efficient strategy for propagation using a dual approach.
-This approach entails loading two frames, specifically target frame $I_{d,t_{1}}$ and reference frame $I_{d,t_{2}}$ (where $I_{d,t_{2}}$ can be randomly selected from the temporal neighbors of $I_{d,t_{1}}$ during training, and is set as the closest frame of $I_{d,t_{1}}$ during inference), propagating features at the deepest layer of the network through our cross-frame interaction (as depicted in Fig.
+This approach entails loading two frames, specifically target frame $\boldsymbol{I}_{d,t_{1}}$ and reference frame $\boldsymbol{I}_{d,t_{2}}$ (where $\boldsymbol{I}_{d,t_{2}}$ can be randomly selected from the temporal neighbors of $\boldsymbol{I}_{d,t_{1}}$ during training, and is set as the closest frame of $\boldsymbol{I}_{d,t_{1}}$ during inference), propagating features at the deepest layer of the network through our cross-frame interaction (as depicted in Fig.
 3), and concurrently supervising these dual outputs to guarantee consistency.
-Figure 4: Visual samples: derive the Retinex decomposition results of enhanced normal-light outputs from the input low-light video data on different datasets.
-Feature propagation via long-range attention and short-range fusion in CFIM.
-Suppose the feature of $I_{d,t_{1}}$ and $I_{d,t_{2}}$ are denoted as $f_{d,t_{1}}$ and $f_{d,t_{2}}$ which is extracted from the same encoder $E$ in the network $f$.
+Feature Propagation via Spatial-varying Fusion.
+Suppose the feature of $\boldsymbol{I}_{d,t_{1}}$ and $\boldsymbol{I}_{d,t_{2}}$ are denoted as $\boldsymbol{f}_{d,t_{1}}$ and $\boldsymbol{f}_{d,t_{2}}$ which is extracted from the same encoder $M_{E}$ in the network $f$.
 To complete the propagation, we initially employ a long-range cross-frame attention operation in the feature space.
-A traditional attention operation Vaswani et al.
-(2017) typically consists of the query vector $q$, the key vector $k$, and the value vector $v$.
-The attention relationship is established using $A(q,k,v)=softmax(q\times k^{T})\times v$, where $\times$ represents matrix multiplication.
+A traditional attention operation (Vaswani et al., 2017) typically consists of the query vector $\boldsymbol{Q}$, the key vector $\boldsymbol{K}$, and the value vector $\boldsymbol{V}$.
+The attention relationship is established using $A(\boldsymbol{Q},\boldsymbol{K},\boldsymbol{V})=\text{softmax}(\boldsymbol{Q}\times\boldsymbol{K}^{\top})\times\boldsymbol{V}$, where $\times$ represents matrix multiplication.
 To enable cross-frame attention, we process the feature via dual paths, as
-$\hat{f}_{d,t_{1}}=A(f_{d,t_{1}},f_{d,t_{1}},f_{d,t_{1}})+A(f_{d,t_{1}},f_{d,t_{2}},f_{d,t_{2}}),$ (9)
-$\hat{f}_{d,t_{2}}=A(f_{d,t_{2}},f_{d,t_{2}},f_{d,t_{2}})+A(f_{d,t_{2}},f_{d,t_{1}},f_{d,t_{1}}).$
+$\boldsymbol{f}^{\prime}_{d,t_{1}}=A(\boldsymbol{f}_{d,t_{1}},\boldsymbol{f}_{d,t_{1}},\boldsymbol{f}_{d,t_{1}}),\boldsymbol{f}^{\prime\prime}_{d,t_{1}}=A(\boldsymbol{f}_{d,t_{1}},\boldsymbol{f}_{d,t_{2}},\boldsymbol{f}_{d,t_{2}}),$ (9)
+$\boldsymbol{f}^{\prime}_{d,t_{2}}=A(\boldsymbol{f}_{d,t_{2}},\boldsymbol{f}_{d,t_{2}},\boldsymbol{f}_{d,t_{2}}),\boldsymbol{f}^{\prime\prime}_{d,t_{2}}=A(\boldsymbol{f}_{d,t_{2}},\boldsymbol{f}_{d,t_{1}},\boldsymbol{f}_{d,t_{1}}).$
 Moreover, a short-range fusion operation is set as the refinement operation to the propagation at both spatial and channel levels. This can be written as
-$\small\bar{f}_{d,t_{1}},\bar{f}_{d,t_{2}}=S(C(\hat{f}_{d,t_{1}}\oplus\hat{f}_{d,t_{2}})),$ (10)
+$(\displaystyle\hat{\boldsymbol{f}}_{d,t_{1}},\hat{\boldsymbol{f}}_{d,t_{2}}=S(C($$(\boldsymbol{f}_{d,t_{1}}+\boldsymbol{f}^{\prime}_{d,t_{1}}+\boldsymbol{f}^{\prime\prime}_{d,t_{1}})\oplus$ (10)
+$(\boldsymbol{f}_{d,t_{2}}+\boldsymbol{f}^{\prime}_{d,t_{2}}+\boldsymbol{f}^{\prime\prime}_{d,t_{2}}))),$
 where $\oplus$ represents channel concatenation, $C$ refers to the convolution network, and $S$ signifies channel dimension splitting.
-The decomposition results for $I_{d,t_{1}}$ and $I_{d,t_{2}}$ are obtained by processing $\bar{f}_{d,t_{1}}$ and $\bar{f}_{d,t_{2}}$ through the decoder.
-We have confirmed that our dual network with synchronous supervision for the outputs of $I_{d,t_{1}}$ and $I_{d,t_{2}}$ can yield superior decomposition and enhancement results compared to individual enhancement strategies.
-3.4 Overall Loss Function
-During the training, we find it is better to adopt the dual training strategy, i.e., for each $I_{d,t_{1}}$, we sample its neighboring reference $I_{d,t_{2}}$ and constrain their outputs simultaneously.
+The decomposition results for $\boldsymbol{I}_{d,t_{1}}$ and $\boldsymbol{I}_{d,t_{2}}$ are obtained by processing $\hat{\boldsymbol{f}}_{d,t_{1}}$ and $\hat{\boldsymbol{f}}_{d,t_{2}}$ through the decoder $M_{D}$.
+We have confirmed that our dual network with synchronous supervision for the outputs of $\boldsymbol{I}_{d,t_{1}}$ and $\boldsymbol{I}_{d,t_{2}}$ can yield superior decomposition and enhancement results compared to individual enhancement strategies.
+The Role of the Cross-attention in CFIM.
+The cross-attention mechanism facilitates information propagation across different frames, aligning with our training procedure that requires spatial consistency for $\boldsymbol{R}$.
+Moreover, during training, CFIM brings varying filtered features from randomly selected reference images ($\boldsymbol{I}_{d,t_{2}}$) into the backbone (i.e., the part without feature propagation), where the input is the image at the current time step $\boldsymbol{I}_{d,t_{1}}$.
+This setup enables CFIM to guide the backbone in learning how to adaptively leverage diverse knowledge from reference images, resulting in robust and effective feature spaces within the backbone, without the need for expensive temporal alignment.
+In other words, the backbone performs strongly without a reference image during inference (though, of course, having a suitable reference image to provide additional useful information generally improves performance), as confirmed by the ablation study.
+Additionally, CFIM can leverage useful cross-frame information when it complements the features in the backbone, further enhancing the decomposition process.
+3.4 Overall Objective to Compute Loss Function
+During the training, we find it is better to adopt the dual training strategy, i.e., for each $\boldsymbol{I}_{d,t_{1}}$, we sample its neighboring reference $\boldsymbol{I}_{d,t_{2}}$ and constrain their outputs simultaneously.
 Thus, the loss function can be written as
-$\small\mathcal{L}_{v}=\mathbb{E}_{t_{1},t_{2}}[\mathcal{L}_{R,t_{1}}+\mathcal{L}_{R,t_{2}}+\lambda_{1}(\mathcal{L}(L_{t_{1}})+\mathcal{L}(L_{t_{2}}))+\lambda_{2}\mathcal{L}(R_{t_{1},t_{2}})],$ (11)
-where $\lambda_{1}$ and $\lambda_{2}$ are the loss weights, and each loss term is defined in Eqs.
-5, 6 and 7.
-Moreover, we observe that the performance of our framework is robust to the choice of the reference frame.
+$\mathcal{O}_{v}=\mathbb{E}_{t_{1},t_{2}}[$ $\mathcal{O}_{rec,t_{1}}+\mathcal{O}_{rec,t_{2}}+$ (11)
+$\lambda_{1}(\mathcal{O}(\boldsymbol{L}_{t_{1}})+\mathcal{O}(\boldsymbol{L}_{t_{2}}))+\lambda_{2}\mathcal{O}(\boldsymbol{R}_{t_{1},t_{2}})],$
+where $\lambda_{1}$ and $\lambda_{2}$ are the loss weights, and each loss term is defined in Equation 5, Equation 6 and Equation 7.
 4 Experiments
+SDSD-indoor SDSD-outdoor SMID DID
+Methods PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM
+SNR 27.30 0.84 23.23 0.82 28.49 0.81 24.85 0.90
+SMG 27.82 0.86 25.17 0.81 28.03 0.78 25.14 0.85
+PairLLE 23.48 0.71 20.04 0.65 22.70 0.63 22.56 0.82
+RetinexFormer 26.56 0.79 22.80 0.77 29.15 0.82 25.40 0.89
+MBLLEN 22.17 0.66 21.41 0.63 22.67 0.68 24.22 0.86
+SMID 24.84 0.72 23.30 0.67 24.78 0.72 22.28 0.84
+SMOID 24.63 0.70 22.25 0.68 23.64 0.71 22.13 0.85
+SDSDNet 26.81 0.75 23.08 0.71 26.03 0.75 22.52 0.81
+DP3DF 27.63 0.74 23.85 0.73 27.19 0.76 22.39 0.88
+StableLLVE 25.32 0.70 22.47 0.65 24.37 0.69 21.64 0.80
+LLVE-SEG 26.19 0.77 24.09 0.72 25.31 0.74 23.85 0.82
+LLVE-CFA 24.28 0.81 22.64 0.76 24.72 0.70 22.53 0.87
+BLLRVE 26.02 0.74 23.50 0.71 26.15 0.73 23.25 0.81
+VRGSS 27.81 0.76 24.28 0.75 27.84 0.78 24.51 0.87
+ZRLLE-PQP 25.53 0.71 22.42 0.73 26.36 0.77 22.84 0.80
+JDAE 26.19 0.74 23.37 0.76 26.45 0.76 23.53 0.82
+Ours 28.93 0.88 26.32 0.82 29.60 0.82 30.10 0.93
+Table 1: Quantitative comparison on SDSD, SMID, and DID datasets. Our method performs the best consistently.
 4.1 Datasets
-Our evaluation is conducted on four publicly available datasets, which encompass a wide range of real-world videos with diverse motion patterns and degradations.
-(1) SMID Chen et al.
-(2019): This dataset consists of static videos with ground truths obtained through long exposure.
-We followed the procedure provided by SMID, converting the videos from the RAW domain to the sRGB domain using rawpy’s default ISP.
-(2) SDSD Wang et al.
-(2021a): SDSD is a dynamic video dataset collected through electromechanical equipment, including both indoor and outdoor subsets.
-(3) DID Fu et al.
-(2023a): DID is another dynamic video dataset characterized by pronounced camera motion, strict spatial alignment, and diverse scene content.
-In comparison to SDSD, DID offers more data and is a multi-illuminance, multi-camera LLVE dataset.
-We follow DID’s division into a training set, a test set, and a validation set, maintaining a ratio of 3:1:1.
-(4) DAVIS Pont-Tuset et al.
-(2017): DAVIS is a dataset that includes both scene dynamics and camera motions.
-To simulate low-light degradation in DAVIS, we followed the methodology outlined in Zhang et al.
-(2021), further incorporating noise degradation.
+Our evaluation is conducted on four publicly available datasets, which encompass a wide range of real-world videos with diverse motion patterns and degradations, including SMID (Chen et al., 2019), SDSD (Wang et al., 2021), DID (Fu et al., 2023a), and DAVIS (Pont-Tuset et al., 2017).
 4.2 Implementation Details
+Experimental Details.
 We conducted experiments on all datasets using the same network structure.
+$T$ is set as 5 in the experiment.
 All modules were trained end-to-end, with the learning rate initialized at $4e^{-4}$ for all layers, adapted by a cosine learning scheduler.
 The batch size used was 4.
-Correspondences were computed using the SOTA method DKM Edstedt et al.
-(2023).
-Additionally, no data augmentation was applied during training.
-All experiments were executed on a single NVIDIA A100 GPU.
-Table 1: Quantitative comparison on SDSD, SMID, and DID datasets. Our method performs the best consistently.
-SDSD-indoor SDSD-outdoor SMID DID
-Methods PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM
-SNR Xu et al. (2022) 27.30 0.84 23.23 0.82 28.49 0.81 24.85 0.90
-SMG Xu et al. (2023b) 27.82 0.86 25.17 0.81 28.03 0.78 25.14 0.85
-PairLLE Fu et al. (2023b) 23.48 0.71 20.04 0.65 22.70 0.63 22.56 0.82
-RetinexFormer Cai et al. (2023) 26.56 0.79 22.80 0.77 29.15 0.82 25.40 0.89
-MBLLEN Lv et al. (2018) 22.17 0.66 21.41 0.63 22.67 0.68 24.22 0.86
-SMID Chen et al. (2019) 24.84 0.72 23.30 0.67 24.78 0.72 22.28 0.84
-SMOID Jiang and Zheng (2019) 24.63 0.70 22.25 0.68 23.64 0.71 22.13 0.85
-SDSDNet Wang et al. (2021a) 26.81 0.75 23.08 0.71 26.03 0.75 22.52 0.81
-DP3DF Xu et al. (2023a) 27.63 0.74 23.85 0.73 27.19 0.76 22.39 0.88
-StableLLVE Zhang et al. (2021) 25.32 0.70 22.47 0.65 24.37 0.69 21.64 0.80
-LLVE-SEG Liu et al. (2023) 26.19 0.77 24.09 0.72 25.31 0.74 23.85 0.82
-LLVE-CFA Chhirolya et al. (2022) 24.28 0.81 22.64 0.76 24.72 0.70 22.53 0.87
-Ours 28.93 0.88 26.32 0.82 29.60 0.82 30.10 0.93
-Table 2: Quantitative comparison on the DAVIS dataset.
-Methods SNR SMG PairLLE RetinexFormer MBLLEN SMID
+Correspondences were computed using the SOTA method DKM (Edstedt et al., 2023).
+We used the pre-trained weights of DKM during training because it is designed for general indoor and outdoor scenes.
+Its generalization ability is supported by its extensive training data and state-of-the-art training strategy, as demonstrated in the original DKM paper and subsequent studies (Zhu and Liu, 2023; Edstedt et al., 2024) (through evaluations on unseen real-world indoor and outdoor scenes).
+4.3 Comparison and Baselines
+We conducted a comprehensive comparison with SOTA LLVE methods, including MBLLEN (Lv et al., 2018), SMID (Chen et al., 2019), SMOID (Jiang and Zheng, 2019), SDSDNet (Wang et al., 2021), DP3DF (Xu et al., 2023a), StableLLVE (Zhang et al., 2021), LLVE-SEG (Liu et al., 2023), LLVE-CFA (Chhirolya et al., 2022), BLLRVE (Zhang et al., 2024), VRGSS (Li et al., 2023), ZRLLE-PQP (Wang et al., 2024), and JDAE (Shi et al., 2024).
+We also compared our method with SOTA LLIE methods for individual frames, including SNR (Xu et al., 2022), SMG (Xu et al., 2023b), PairLLE (Fu et al., 2023b), RetinexFormer (Cai et al., 2023).
+All these methods were trained on each dataset using their respective released code and hyper-parameters (e.g., the training epoch that is set to guarantee convergence).
+Note that all baselines are trained on our unified data split for a fair comparison.
+The splits slightly differ from those in baselines’ original papers, so the scores of baselines may vary from the original papers.
+Input
+RetinexFormer
+SDSDNet
+DP3DF
+Ours
+GT
+Figure 4: Visual comparisons on SDSD-indoor (top two rows), SDSD-outdoor (middle two rows), and SMID (bottom two rows).
+The results of our proposed framework, i.e., “Ours”, demonstrate better visual perception with clearer visibility and more enhanced details.
+Input
+RetinexFormer
+SDSDNet
+DP3DF
+Ours
+GT
+Figure 5: Comparisons on DID (top two rows) and DAVIS (bottom two rows). “Ours” has better visibility and details.
+SNR SMG PairLLE RetinexFormer MBLLEN SMID
 PSNR 20.69 20.18 19.57 21.79 18.63 20.51
 SSIM 0.710 0.672 0.667 0.723 0.619 0.686
-Methods SMOID SDSDNet DP3DF StableLLVE LLVE-SEG Ours
+SMOID SDSDNet DP3DF StableLLVE LLVE-SEG Ours
 PSNR 20.66 21.22 22.04 21.48 21.45 23.38
 SSIM 0.697 0.716 0.749 0.732 0.715 0.782
-4.3 Comparison and Baselines
-We conducted a comprehensive comparison with SOTA LLVE methods, including MBLLEN Lv et al.
-(2018), SMID Chen et al.
-(2019), SMOID Jiang and Zheng (2019), SDSDNet Wang et al.
-(2021a), DP3DF Xu et al.
-(2023a), StableLLVE Zhang et al.
-(2021), LLVE-SEG Liu et al.
-(2023), and LLVE-CFA Chhirolya et al.
-(2022) We also compared our method with SOTA LLIE methods for individual frames, including SNR Xu et al.
-(2022), SMG Xu et al.
-(2023b), PairLLE Fu et al.
-(2023b), RetinexFormer Cai et al.
-(2023).
-All these methods were trained on each dataset using their respective released code and hyper-parameters (e.g., the training epoch that is set to guarantee the convergence).
-Note that all these baselines are trained on our unified data split for a fair comparison.
-The splits slightly differ from those in baselines’ original papers, so the scores of the baselines may vary from the original papers.
+Table 2: Quantitative comparison on the DAVIS dataset.
 Quantitative Result.
@@
 Furthermore, Table 2 provides a summary of the comparative results on DAVIS.
-In comparison to the degradation synthesis strategy as presented in Zhang et al.
-(2021), we have extended our approach to include the degradation of a low-light noise term.
+In comparison to the degradation synthesis strategy as presented in (Zhang et al., 2021), we have extended our approach to include the degradation of a low-light noise term.
 As a result, our synthesized data encompasses dynamic scenes with pronounced invisibility and perturbations, posing a considerable challenge.
 As illustrated in Table 2, our method consistently yields the highest PSNR and SSIM values, reaffirming the effectiveness of our approach.
+DAVIS SDSD-Indoor SDSD-Outdoor DID
+Methods Short Long Short Long Short Long Short Long
+SNR 0.027 0.070 0.017 0.060 0.022 0.046 0.025 0.068
+RetinexFormer 0.029 0.072 0.018 0.061 0.024 0.043 0.024 0.065
+SDSDNet 0.024 0.068 0.012 0.053 0.011 0.038 0.017 0.059
+DP3DF 0.021 0.065 0.011 0.050 0.013 0.036 0.019 0.062
+StableLLVE 0.023 0.067 0.016 0.055 0.019 0.042 0.023 0.066
+LLVE-SEG 0.025 0.071 0.014 0.057 0.016 0.040 0.021 0.064
+Ours-R 0.017 0.058 0.010 0.048 0.007 0.030 0.011 0.049
+Ours 0.020 0.063 0.012 0.051 0.010 0.034 0.015 0.054
+Table 3: The quantitative comparison in terms of short-term (“Short”) and long-term (“Long”) temporal loss. “Ours-R” means the view-independent term produced by ours.
+Evaluation for Temporal Consistency.
+For video processing, the performance of temporal consistency and stability should be evaluated.
+Thus, we employ the short-term and long-term temporal loss proposed in (Lai et al., 2018) for such temporal evaluation on different datasets.
+The wrapping operations among frames are computed on the frames with the normal light.
+Moreover, the long-term loss is computed every 10 frames.
+The results are shown in Table 3 (we normalize the frame values into [0, 1]).
+Obviously, our results have lower temporal loss than the baselines, e.g., higher temporal consistency and stability.
+Specifically, we assess the temporal consistency of the view-independent term.
+The results presented in Table 3 further validate the constancy.
 Qualitative Result.
@@
 Fig.
-5 showcases the visual comparisons on SDSD and SMID, while Fig.
+4 showcases the visual comparisons of SDSD and SMID, while Fig.
 5 displays visual cases from DID and DAVIS.
-In general, the results enhanced by our approach exhibit a more natural appearance, including accurate color, well-balanced brightness, enhanced contrast, and precise details without noise.
-Furthermore, our framework’s results show fewer artifacts in regions with complex textures, and they appear cleaner and sharper compared to the results produced by other methods.
+In general, the results enhanced by our approach exhibit a more natural appearance, including accurate color, well-balanced brightness, enhanced contrast, and precise details.
+Our results show fewer artifacts in regions with complex textures, and they appear cleaner and sharper compared to results produced by other methods.
 This distinction is particularly notable, as most baselines perform well in simpler areas.
-Input
-RetinexFormer
-SDSDNet
-DP3DF
-Ours
-GT
-Input
-RetinexFormer
-SDSDNet
-DP3DF
-Ours
-GT
-Input
-RetinexFormer
-SDSDNet
-DP3DF
-Ours
-GT
-Figure 5: Visual comparisons on SDSD-indoor (top), SDSD-outdoor (middle), and SMID (bottom).
-The results of our proposed framework, i.e., “Ours”, demonstrate better visual perception with less noise, clearer visibility, and more enhanced details.
-Input
-RetinexFormer
-SDSDNet
-DP3DF
-Ours
-GT
-Input
-RetinexFormer
-SDSDNet
-DP3DF
-Ours
-GT
-Figure 6: Comparisons on DID (top) and DAVIS (bottom). “Ours” has less noise, better visibility and details.
-Figure 7: The above pie charts summarize the results of our user study, and ours is preferred by participants.
-4.4 Ablation Study
-We assess the critical components of our framework through five ablation cases.
-(1) “w/o LR” that removes the constraint on learning the reflectance part.
-(2) “w/o LL” where the constraint on learning the illumination part is removed.
-(3) “w/o C.F.”: we eliminate the cross-frame attention and fusion in the network, resulting in each frame being enhanced individually.
-(4) “with M.I.”: This refers to using multiple neighboring frames as input and employing the temporal alignment of deformable convolution.
-(5) “w/o Dual”: the dual learning strategy is removed, meaning that the loss is applied to only one frame in each iteration.
-The results are summarized in Table 3.
-By comparing “w/o LR” with the full setting (“Full”), we can clearly demonstrate the effectiveness of our proposed constraint for the reflectance map.
-Similarly, the significant advantage of “Full” over “w/o LL” highlights the significance of the proposed illumination continuity constraint.
-Furthermore, we can validate the importance of mutually propagating features among different frames to achieve spatial-temporal consistent decomposition by comparing “w/o C.F.” with “Full”.
-Our proposed propagation strategy via the dual network structure proves to be more effective than the common method of using multi-frame inputs for propagation in videos, as seen in the comparison between “w/o M.I.” and “Full”.
-The dual learning setting, where we simultaneously supervise the two outputs of the dual network, also proves to be valuable, as evidenced by the comparison between “w/o Dual” and “Full”.
-Table 3: Ablation study on SDSD, SMID, and DID.
 SDSD-indoor SDSD-outdoor SMID DID
@@
 Full 28.93 0.88 26.32 0.82 29.60 0.82 30.10 0.93
+Table 4: Ablation study on SDSD, SMID, and DID.
+Figure 6: The above pie charts summarize the results of our user study, and ours is preferred by participants.
+4.4 Ablation Study
+We assess the critical components of our framework through five ablation cases.
+(1) “w/o LR” that removes the constraint on learning the view-independent part.
+(2) “w/o LL” where the constraint on learning the view-dependent part is removed.
+(3) “w/o C.F.”: we eliminate the cross-frame attention and fusion in the network, resulting in each frame being enhanced individually.
+(4) “with M.I.”: This refers to using multiple neighboring frames as input and employing the temporal alignment of deformable convolution.
+(5) “w/o Dual”: the dual learning strategy is removed, meaning that the loss is applied to only one frame in each iteration.
+The results are summarized in Table 4.
+By comparing “w/o LR” with the full setting (“Full”), we can clearly demonstrate the effectiveness of our proposed constraint for the view-independent part.
+Similarly, the significant advantage of “Full” over “w/o LL” highlights the significance of the proposed continuity constraint.
+Furthermore, we can validate the importance of mutually propagating features among different frames to achieve spatial-temporal consistent decomposition by comparing “w/o C.F.” with “Full”.
+Our proposed propagation strategy via the dual network structure proves to be more effective than the common method of using multi-frame inputs for propagation in videos, as seen in the comparison between “w/o M.I.” and “Full”.
+The dual learning setting, where we simultaneously supervise the two outputs of the dual network, also proves to be valuable, as evidenced by the comparison between “w/o Dual” and “Full”.
 4.5 User Study
 To prove the effectiveness of our proposed framework in terms of human subjective evaluation, we conduct a large-scale user study with 100 participants with varying ages and education backgrounds, and balanced sex distribution.
-We aim to compare the perceptual quality of our method against SOTA baselines.
-To conduct the user study, we randomly collect 30 videos in the wild with the capture device of Sony A7R3 and iPhone13 (all models in the user study are trained with DID, since it is a large-scale dataset with multiple scenes and devices).
-Following most of the existing low-light enhancement works Xu et al.
-(2023b); Wang et al.
-(2023a), we employ the AB-test for the user study.
-For a video, the result produced by our method is called “Video A” whereas the result from another baseline is “Video B”.
-During the evaluation, each participant will compare the results of our method against other baselines in random order.
-For the comparison with one baseline, 5 videos will be randomly selected for visual perception.
-Each participant would simultaneously watch videos A and B.
-To avoid the bias caused by the position, we will randomize the left-right presentation order when showing videos A and B in each question.
-The participants can choose among three options: “Video A is better”, “Video B is better”, and “I think they are of the same quality”.
+Following most of the existing low-light enhancement works (Xu et al., 2023b; Wang et al., 2023a), we employ the AB-test for the user study.
+Participants should indicate their preference or select the same option.
 They should make the decision according to the natural brightness, contrast, and color of each frame; the rich details, fewer artifacts, and the temporal consistency in the video.
-For each participant, the number of tasks is 10 methods $\times$ 5 videos $=50$, and it took around 50 minutes on average for each participant to complete the user study.
 Fig.
@@
 This demonstrates that our method’s results are more preferred by the human subjective perception.
-Table 4: Quantitative comparison on 3D low-light datasets with normal-light pairs, proposed in Wang et al. (2023a).
-Methods PairLLE SNR RetinexFormer SCI+NeRF Ours
-PSNR 18.35 18.87 19.49 13.08 21.09
-SSIM 0.604 0.650 0.702 0.505 0.763
-Methods PairLLE+NeRF SNR+NeRF RetinexFormer+NeRF LLNeRF Ours+NeRF
-PSNR 18.72 19.24 19.98 20.50 22.35
-SSIM 0.681 0.693 0.736 0.758 0.782
-4.6 Enhancement for NeRF
-Recent works have introduced the implementation of 3D low-light enhancement Wang et al.
-(2023a), with a primary focus on disentangling view-independent intrinsic color and view-varying illuminations to achieve 3D-consistent enhancement.
-On the other hand, our framework’s enhanced results are particularly well-suited for NeRF reconstruction, as we have implemented a decomposition that maintains consistency across different spatial locations within a 3D scene.
-To assess the impact of using our enhanced results for NeRF, we conducted experiments on a multi-view low-light dataset previously introduced in Wang et al.
-(2023a).
-We employed the network trained on DID, as it benefits from a large-scale training dataset, resulting in solid generalization capabilities.
-Furthermore, DID inherently contains multi-view images.
-Specifically, for each view, we selected its closest view as the reference view for our input data.
-In Table 4, we present the PSNR and SSIM results on the testing set of Wang et al.
-(2023a), where ground truth data is available.
-It’s important to note that none of the networks were fine-tuned on the testing set.
-The superiority of our framework is evident, as our network can achieve 3D-consistent enhancement that aligns with the fundamental principles of NeRF.
+4.6 The Effects of CFIM Structure
+As previously discussed, the use of CFIM yields two key benefits: it enhances the consistency of the image enhancement process and strengthens the robustness of the backbone by providing the cross-attention mechanism with information from random neighboring time steps.
+In this section, we conduct experiments to show the effects of CFIM.
+The baseline consists of a network that processes a single frame, while the comparative setup integrates CFIM, where cross-attention is applied between randomly selected pairs of frames during training.
+The results of this comparison are presented in Table 5, which show that models incorporating the CFIM structure significantly outperform those without it, thereby demonstrating the effectiveness of the CFIM approach.
+Besides, the ablation results of the mentioned “w/o C.F.” (in Sec.
+4.4) also support the effects of CFIM in the decomposition.
+SDSD-indoor SDSD-outdoor SMID DID
+Methods PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM
+SNR 27.30 0.84 23.23 0.82 28.49 0.81 24.85 0.90
+SNR+CFIM 28.15 0.86 24.57 0.83 29.24 0.82 26.03 0.91
+RetinexFormer 26.56 0.79 22.80 0.77 29.15 0.82 25.40 0.89
+RetinexFormer+CFIM 27.42 0.80 23.75 0.79 29.87 0.83 26.38 0.90
+Table 5: The experiments to show the effects of CFIM.
+4.7 Robustness Towards Correspondences
+In this section, we explore the robustness of our framework with respect to incorrect and insufficient correspondences.
+In the first experiment, we introduce random noise perturbations to the computed correspondences, modifying the location values of the correspondences (the perturbation range is sampled from $-20\sim 20$ pixels, which falls within the normal error range of current correspondence estimation methods (Edstedt et al., 2023)).
+As shown in Table 6, the results with perturbed correspondences (denoted as Ours with per.) show a performance drop compared to the unperturbed case.
+However, the performance remains superior to that of most SOTA baselines, indicating that our framework is robust to a certain of erroneous correspondences, which are likely to occur in areas with complex textures or other challenging scenarios.
+In the second experiment, we address the issue of insufficient correspondences, which can arise, for instance, in regions with occlusions.
+To simulate this, we randomly reduce the number of correspondences to 10% of the originally detected ones.
+The results (Ours with reduced) are presented in Table 6, where we observe that our method continues to outperform almost all baselines in Table 1, albeit with a smaller margin than the original setting.
+This suggests that even with fewer correspondences, the remaining ones still provide valuable information and they are in accord with the other constraints.
+This further demonstrates the robustness of our method.
+SDSD-indoor SDSD-outdoor SMID DID
+Methods PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM
+Ours with reduce 27.97 0.84 25.73 0.80 29.12 0.79 28.86 0.89
+Ours with per. 28.10 0.86 26.04 0.79 29.37 0.81 29.08 0.91
+Ours 28.93 0.88 26.32 0.82 29.60 0.82 30.10 0.93
+Table 6: The experiments to show the robustness of our framework towards correspondences.
 5 Conclusion
-We introduce a novel LLVE framework that utilizes spatial-temporal consistent illumination and reflection decomposition and dual networks for mutual feature propagation.
-It predicts the decomposition of normal-light outputs without external priors, thanks to cross-frame correspondences and practical illumination continuity.
+We introduce a novel LLVE framework that utilizes spatial-temporal consistent decomposition and dual networks for mutual feature propagation.
+It predicts the decomposition of normal-light outputs without external priors, thanks to cross-frame correspondences and practical constraint of the continuity.
 Extensive experiments on various datasets showcase the framework’s superiority over SOTA approaches and the impact of our designed components.
+Ethical Statement
+There are no ethical issues.
+Acknowledgments
+This work is supported by the Natural Science Foundation of Zhejiang Province, China, under No. LD24F020002.

### v3

@@
 1 Introduction
-Low-light enhancement aims to enhance underexposed images and videos captured in low-light conditions (Xu et al., 2022; Wang et al., 2021), improving their visual quality while reducing noise.
-This technique can be applied in wide-ranging applications, such as portrait photography on mobile devices (Ignatov et al., 2017; Hasinoff et al., 2016), nighttime face recognition (Ma et al., 2022; Wang et al., 2022) and vehicle detection (Fu et al., 2023a; Wu et al., 2022).
+Low-light enhancement aims to enhance underexposed images and videos captured in low-light conditions Xu et al.
+(2022); Wang et al.
+(2021), improving their visual quality while reducing noise.
+This technique can be applied in wide-ranging applications, such as portrait photography on mobile devices Ignatov et al.
+(2017); Hasinoff et al.
+(2016), nighttime face recognition Ma et al.
+(2022); Wang et al.
+(2022) and vehicle detection Fu et al.
+(2023a); Wu et al.
+(2022).
 The key challenge when enhancing videos is the need for consistent enhancement results across corresponding locations in different frames, which vary both spatially and temporally.
@@
 One solution is to incorporate parameterized enhancement models with physical significance into the enhancement process, reducing the reliance on training data.
-According to the linear Lambertian model for intrinsic image decomposition (Yuille, 2012; Narihira et al., 2015), an observed image $\boldsymbol{I}$ can be expressed as the element-wise product ($\otimes$) of its albedo component and shading component.
+According to the linear Lambertian model for intrinsic image decomposition Yuille (2012); Narihira et al.
+(2015), an observed image $\boldsymbol{I}$ can be expressed as the element-wise product ($\otimes$) of its albedo component and shading component.
 The albedo part is normally view-independent, capturing the intrinsic appearance, whereas the shading component is commonly view-dependent, influenced by lighting conditions and the surface normal’s direction.
@@
 Two frames are used as inputs, and their features are propagated within the designed Cross-Frame Interaction Module (CFIM).
-CFIM differs from other similar architectures in the restoration task (Lv et al., 2023) in terms of the fusion manner.
+CFIM differs from other similar architectures in the restoration task Lv et al.
+(2023) in terms of the fusion manner.
 CFIM facilitates interaction through a combination of long-range cross-frame attention computation and short-range channel-spatial fusion, ensuring the accurate and comprehensive merging of features from both global and local perspectives.
@@
 To enhance the quality of a low-light video, image enhancement methods can be applied on a frame-by-frame basis.
-In recent years, learning-based Low-Light Image Enhancement (LLIE) techniques (Jiang et al., 2021; Yang et al., 2021) have made significant advancements, with a primary focus on supervised approaches.
+In recent years, learning-based Low-Light Image Enhancement (LLIE) techniques Jiang et al.
+(2021); Yang et al.
+(2021) have made significant advancements, with a primary focus on supervised approaches.
 Low-Light Video Enhancement.
 In addition to the need for LLIE, there is a growing demand for video enhancement, considering the widespread use of videos as a popular data format on the internet and in photographic equipment.
-Various approaches have been proposed in this context (Chen et al., 2019; Triantafyllidou et al., 2020; Ye et al., 2023; Lv et al., 2023; Xu et al., 2023a; Fu et al., 2023a).
+Various approaches have been proposed in this context Chen et al.
+(2019); Triantafyllidou et al.
+(2020); Ye et al.
+(2023); Lv et al.
+(2023); Xu et al.
+(2023a); Fu et al.
+(2023a).
 Liu et.al.
-(Liu et al., 2023) and Liang et.al.
-(Liang et al., 2023) used prior event information to learn enhancement mapping for brightening videos.
+Liu et al.
+(2023) and Liang et.al.
+Liang et al.
+(2023) used prior event information to learn enhancement mapping for brightening videos.
 Xu et.al.
-(Xu et al., 2023a) designed a parametric 3D filter tailored for enhancing and sharpening low-light videos.
+Xu et al.
+(2023a) designed a parametric 3D filter tailored for enhancing and sharpening low-light videos.
 Recently, Fu et.al.
-(Fu et al., 2023a) introduced a video enhancement method called LAN, which iteratively refines illumination and adaptively adjusts it.
+Fu et al.
+(2023a) introduced a video enhancement method called LAN, which iteratively refines illumination and adaptively adjusts it.
 However, it’s important to note that LAN lacks an explicit constraint for maintaining consistent reflectance and illumination decomposition.
-Several datasets for video enhancement, encompassing both static (Chen et al., 2019; Jiang and Zheng, 2019; Wang et al., 2019; Triantafyllidou et al., 2020) and dynamic motions (Wang et al., 2021; Fu et al., 2023a), have been introduced.
+Several datasets for video enhancement, encompassing both static Chen et al.
+(2019); Jiang and Zheng (2019); Wang et al.
+(2019); Triantafyllidou et al.
+(2020) and dynamic motions Wang et al.
+(2021); Fu et al.
+(2023a), have been introduced.
 In this paper, we introduce a novel LLVE method, designed to enhance effects on these datasets.
@@
 Motivation.
-According to the linear Lambertian model (Narihira et al., 2015), an observed image $\boldsymbol{I}$ can be formulated as the element-wise product of its albedo and shading component.
+According to the linear Lambertian model Narihira et al.
+(2015), an observed image $\boldsymbol{I}$ can be formulated as the element-wise product of its albedo and shading component.
 Here, the albedo part is view-independent, capturing the intrinsic appearance, while the shading component is view-dependent.
@@
 3.2).
-When applying this model to sequential video data, the objective related to the decomposition of video data can be summarized as
+When applying this model to sequential video data, the objective related to the decomposition of video can be written as
 $\mathcal{O}_{v}=\mathbb{E}_{t=1:T}[\mathcal{O}_{rec,t}+\mathcal{O}(\boldsymbol{L}_{t})+\mathcal{O}(\boldsymbol{R}_{t})],$ (3)
@@
 where $\mathbb{E}$ is the average operation, $T$ is the number of frame, and $\boldsymbol{L}_{t}$/$\boldsymbol{R}_{t}$ is the decomposition output at the time index of $t$.
-Compared with Equation 2, we need to simultaneously guarantee the decomposition quality and the consistency among different frames’ results.
+Compared with Equation 2, we need to simultaneously guarantee the decomposition quality and the temporal consistency.
 Video Data Guide the Decomposition by Themselves.
@@
 Motivation.
-Video can be conceptualized as static/dynamic 3D multi-view data (Wang et al., 2023b, a).
-Prior research has demonstrated that the multi-view data itself can be harnessed to decompose the view-dependent and view-independent elements, which respectively encapsulate the intrinsic texture and view-altering illuminations (Wang et al., 2023a).
+Video can be conceptualized as static/dynamic 3D multi-view data Wang et al.
+(2023b, a).
+Prior research has demonstrated that the multi-view data itself can be harnessed to decompose the view-dependent and view-independent elements, which respectively encapsulate the intrinsic texture and view-altering illuminations Wang et al.
+(2023a).
 Thus, once we establish correspondence relationships among $\boldsymbol{I}_{d,t}$ for all $t$, we can apply the view-independent constraint to derive $\hat{\boldsymbol{R}}_{n,t}$, as it represents the intrinsic properties of the target scene.
-Subsequently, obtaining $\hat{\boldsymbol{L}}_{n,t}$ becomes achievable through the utilization of the inherent image reconstruction loss, in conjunction with adherence to the continuity assumption.
+Subsequently, obtaining $\hat{\boldsymbol{L}}_{n,t}$ becomes achievable through utilizing the inherent reconstruction loss, in conjunction with adherence to the continuity assumption.
 Implementation of Correspondences.
@@
 2, where $\boldsymbol{I}_{d,t}$ and $\boldsymbol{I}_{n,t}$ are pixel-wise aligned.
-Given two frames $\boldsymbol{I}_{n,t_{1}}$ and $\boldsymbol{I}_{n,t_{2}}$, correspondences can be determined using the prediction network $F$ (Edstedt et al., 2023).
+Given two frames $\boldsymbol{I}_{n,t_{1}}$ and $\boldsymbol{I}_{n,t_{2}}$, correspondences can be determined using the prediction network $F$ Edstedt et al.
+(2023).
 We assume the detection of $M$ correspondences, denoted as ${c_{m}=(x_{1,m},y_{1,m},x_{2,m},y_{2,m}),m\in[1,M]}$.
@@
 This observation is empirically validated in our ablation study.
-Current LLVE methods primarily leverage multiple-frame inputs to facilitate propagation (Wang et al., 2021; Fu et al., 2023a).
-However, these strategies come with a notable cost in terms of aligning features, potentially affecting inference efficiency.
+Current LLVE methods primarily leverage multiple-frame inputs to facilitate propagation Wang et al.
+(2021); Fu et al.
+(2023a).
+However, these strategies come with a notable cost in terms of aligning features.
 In this paper, we introduce an efficient strategy for propagation using a dual approach.
@@
 To complete the propagation, we initially employ a long-range cross-frame attention operation in the feature space.
-A traditional attention operation (Vaswani et al., 2017) typically consists of the query vector $\boldsymbol{Q}$, the key vector $\boldsymbol{K}$, and the value vector $\boldsymbol{V}$.
+A traditional attention operation Vaswani et al.
+(2017) typically consists of the query vector $\boldsymbol{Q}$, the key vector $\boldsymbol{K}$, and the value vector $\boldsymbol{V}$.
 The attention relationship is established using $A(\boldsymbol{Q},\boldsymbol{K},\boldsymbol{V})=\text{softmax}(\boldsymbol{Q}\times\boldsymbol{K}^{\top})\times\boldsymbol{V}$, where $\times$ represents matrix multiplication.
@@
 This setup enables CFIM to guide the backbone in learning how to adaptively leverage diverse knowledge from reference images, resulting in robust and effective feature spaces within the backbone, without the need for expensive temporal alignment.
-In other words, the backbone performs strongly without a reference image during inference (though, of course, having a suitable reference image to provide additional useful information generally improves performance), as confirmed by the ablation study.
-Additionally, CFIM can leverage useful cross-frame information when it complements the features in the backbone, further enhancing the decomposition process.
+CFIM also leverages cross-frame information when it complements backbone.
 3.4 Overall Objective to Compute Loss Function
@@
 4.1 Datasets
-Our evaluation is conducted on four publicly available datasets, which encompass a wide range of real-world videos with diverse motion patterns and degradations, including SMID (Chen et al., 2019), SDSD (Wang et al., 2021), DID (Fu et al., 2023a), and DAVIS (Pont-Tuset et al., 2017).
+Our evaluation is conducted on four publicly available datasets, which encompass a wide range of real-world videos with diverse motion patterns and degradations, including SMID Chen et al.
+(2019), SDSD Wang et al.
+(2021), DID Fu et al.
+(2023a), and DAVIS Pont-Tuset et al.
+(2017).
 4.2 Implementation Details
@@
 The batch size used was 4.
-Correspondences were computed using the SOTA method DKM (Edstedt et al., 2023).
+Correspondences were computed using the SOTA method DKM Edstedt et al.
+(2023).
 We used the pre-trained weights of DKM during training because it is designed for general indoor and outdoor scenes.
-Its generalization ability is supported by its extensive training data and state-of-the-art training strategy, as demonstrated in the original DKM paper and subsequent studies (Zhu and Liu, 2023; Edstedt et al., 2024) (through evaluations on unseen real-world indoor and outdoor scenes).
+Its generalization ability is supported by its extensive training data and state-of-the-art training strategy, as demonstrated in the original DKM paper and subsequent studies Zhu and Liu (2023); Edstedt et al.
+(2024) (through evaluations on unseen scenes).
 4.3 Comparison and Baselines
-We conducted a comprehensive comparison with SOTA LLVE methods, including MBLLEN (Lv et al., 2018), SMID (Chen et al., 2019), SMOID (Jiang and Zheng, 2019), SDSDNet (Wang et al., 2021), DP3DF (Xu et al., 2023a), StableLLVE (Zhang et al., 2021), LLVE-SEG (Liu et al., 2023), LLVE-CFA (Chhirolya et al., 2022), BLLRVE (Zhang et al., 2024), VRGSS (Li et al., 2023), ZRLLE-PQP (Wang et al., 2024), and JDAE (Shi et al., 2024).
-We also compared our method with SOTA LLIE methods for individual frames, including SNR (Xu et al., 2022), SMG (Xu et al., 2023b), PairLLE (Fu et al., 2023b), RetinexFormer (Cai et al., 2023).
-All these methods were trained on each dataset using their respective released code and hyper-parameters (e.g., the training epoch that is set to guarantee convergence).
+We conducted a comprehensive comparison with SOTA LLVE methods, including MBLLEN Lv et al.
+(2018), SMID Chen et al.
+(2019), SMOID Jiang and Zheng (2019), SDSDNet Wang et al.
+(2021), DP3DF Xu et al.
+(2023a), StableLLVE Zhang et al.
+(2021), LLVE-SEG Liu et al.
+(2023), LLVE-CFA Chhirolya et al.
+(2022), BLLRVE Zhang et al.
+(2024), VRGSS Li et al.
+(2023), ZRLLE-PQP Wang et al.
+(2024), and JDAE Shi et al.
+(2024).
+We also compared our method with SOTA LLIE methods for individual frames, including SNR Xu et al.
+(2022), SMG Xu et al.
+(2023b), PairLLE Fu et al.
+(2023b), RetinexFormer Cai et al.
+(2023).
+All methods were trained on each dataset using their respective released code and hyper-parameters (e.g., the training epoch that is set to guarantee convergence).
 Note that all baselines are trained on our unified data split for a fair comparison.
-The splits slightly differ from those in baselines’ original papers, so the scores of baselines may vary from the original papers.
+The splits slightly differ from those in baselines’ original papers, so scores of baselines may vary from original papers.
 Input
@@
 Furthermore, Table 2 provides a summary of the comparative results on DAVIS.
-In comparison to the degradation synthesis strategy as presented in (Zhang et al., 2021), we have extended our approach to include the degradation of a low-light noise term.
+In comparison to the degradation synthesis strategy as presented in Zhang et al.
+(2021), we have extended our approach to include the degradation of a low-light noise term.
 As a result, our synthesized data encompasses dynamic scenes with pronounced invisibility and perturbations, posing a considerable challenge.
@@
 For video processing, the performance of temporal consistency and stability should be evaluated.
-Thus, we employ the short-term and long-term temporal loss proposed in (Lai et al., 2018) for such temporal evaluation on different datasets.
+Thus, we employ the short-term and long-term temporal loss proposed in Lai et al.
+(2018) for such temporal evaluation on different datasets.
 The wrapping operations among frames are computed on the frames with the normal light.
@@
 To prove the effectiveness of our proposed framework in terms of human subjective evaluation, we conduct a large-scale user study with 100 participants with varying ages and education backgrounds, and balanced sex distribution.
-Following most of the existing low-light enhancement works (Xu et al., 2023b; Wang et al., 2023a), we employ the AB-test for the user study.
+Following most of the existing low-light enhancement works Xu et al.
+(2023b); Wang et al.
+(2023a), we employ the AB-test for the user study.
 Participants should indicate their preference or select the same option.
@@
 SNR+CFIM 28.15 0.86 24.57 0.83 29.24 0.82 26.03 0.91
-RetinexFormer 26.56 0.79 22.80 0.77 29.15 0.82 25.40 0.89
-RetinexFormer+CFIM 27.42 0.80 23.75 0.79 29.87 0.83 26.38 0.90
-Table 5: The experiments to show the effects of CFIM.
+R.F. 26.56 0.79 22.80 0.77 29.15 0.82 25.40 0.89
+R.F.+CFIM 27.42 0.80 23.75 0.79 29.87 0.83 26.38 0.90
+Table 5: The effects of CFIM. “R.F." denotes RetinexFormer.
 4.7 Robustness Towards Correspondences
 In this section, we explore the robustness of our framework with respect to incorrect and insufficient correspondences.
-In the first experiment, we introduce random noise perturbations to the computed correspondences, modifying the location values of the correspondences (the perturbation range is sampled from $-20\sim 20$ pixels, which falls within the normal error range of current correspondence estimation methods (Edstedt et al., 2023)).
-As shown in Table 6, the results with perturbed correspondences (denoted as Ours with per.) show a performance drop compared to the unperturbed case.
+In the first experiment, we introduce random noise perturbations to the computed correspondences, modifying the location values of the correspondences (the perturbation range is sampled from $-20\sim 20$ pixels, which falls within the normal error range of current correspondence estimation methods Edstedt et al.
+(2023)).
+As shown in Table 6, the results with perturbed correspondences (denoted as “Ours with per.") show a performance drop compared to the unperturbed case.
 However, the performance remains superior to that of most SOTA baselines, indicating that our framework is robust to a certain of erroneous correspondences, which are likely to occur in areas with complex textures or other challenging scenarios.
@@
 To simulate this, we randomly reduce the number of correspondences to 10% of the originally detected ones.
-The results (Ours with reduced) are presented in Table 6, where we observe that our method continues to outperform almost all baselines in Table 1, albeit with a smaller margin than the original setting.
+The results (“Ours with red.") are presented in Table 6, where we observe that our method continues to outperform almost all baselines in Table 1, albeit with a smaller margin than the original setting.
 This suggests that even with fewer correspondences, the remaining ones still provide valuable information and they are in accord with the other constraints.
@@
 Methods PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM
-Ours with reduce 27.97 0.84 25.73 0.80 29.12 0.79 28.86 0.89
+Ours with red. 27.97 0.84 25.73 0.80 29.12 0.79 28.86 0.89
 Ours with per. 28.10 0.86 26.04 0.79 29.37 0.81 29.08 0.91
 Ours 28.93 0.88 26.32 0.82 29.60 0.82 30.10 0.93
-Table 6: The experiments to show the robustness of our framework towards correspondences.
+Table 6: The robustness of our framework towards correspondences.
 5 Conclusion